### v1

# MoVieS: Motion-Aware 4D Dynamic View Synthesis in One Second

Abstract
We present MoVieS, a novel feed-forward model that synthesizes 4D dynamic novel views from monocular videos in one second. MoVieS represents dynamic 3D scenes using pixel-aligned grids of Gaussian primitives, explicitly supervising their time-varying motion. This allows, for the first time, the unified modeling of appearance, geometry and motion, and enables view synthesis, reconstruction and 3D point tracking within a single learning-based framework. By bridging novel view synthesis with dynamic geometry reconstruction, MoVieS enables large-scale training on diverse datasets with minimal dependence on task-specific supervision. As a result, it also naturally supports a wide range of zero-shot applications, such as scene flow estimation and moving object segmentation. Extensive experiments validate the effectiveness and efficiency of MoVieS across multiple tasks, achieving competitive performance while offering several orders of magnitude speedups.
[…]
1 Introduction
Humans and animals perceive a continuous stream of observations from a dynamic three-dimensional world, and effortlessly interpret its underlying 3D geometry and motion. Replicating this capability is essential for any embodied agent that must understand and act in the physical world.
Recent advances have made great strides in individual 3D tasks, such as monocular depth estimation (ranftl2020towards, ; ke2024repurposing, ; yang2024depth, ; bochkovskii2024depth, ; chen2025video, ), 3D scene reconstruction (schonberger2016structure, ; teed2021droid, ; wang2024dust3r, ; zhang2024monst3r, ; wang2025vggt, ), novel view synthesis (mildenhall2020nerf, ; fridovich2023k, ; kerbl20233d, ; wu20244d, ; luiten2024dynamic, ) and point tracking (harley2022particle, ; doersch2023tapir, ; wang2023tracking, ; karaev2024cotracker, ; xiao2024spatialtracker, ). However, besides treating each task in isolation, most existing view synthesis and reconstruction studies focus on static scenes and require costly per-scene optimization without learning prior knowledge. Real-world environments are inherently dynamic and diverse, and all the aforementioned 3D scene understanding tasks could share common underlying principles.
Motivated by this, we introduce MoVieS, a Motion-aware dynamic View Synthesis model for feed-forward 4D reconstruction of videos in the wild, that jointly models scenes’ appearance, geometry and motion. MoVieS represents 3D dynamic scenes using renderable and deformable 3D particles, termed dynamic splatter pixels, and utilizes a differentiable 3D Gaussian rendering framework (kerbl20233d, ). Specifically, following recent practices on feed-forward view synthesis (charatan2024pixelsplat, ; szymanowicz2024splatter, ; zhang2024gs, ; xu2024depthsplat, ; lin2025diffsplat, ), each input pixel is mapped to a 3D Gaussian primitive, with its 3D location determined by predicted depth. To model dynamics, MoVieS regresses per-pixel motion displacements toward arbitrary query timestamps, enabling temporal tracking of each splatter pixel. This design facilitates coherent reconstruction of both 3D geometry and appearance across camera viewpoints and temporal frames.
As shown in Figure 1, MoVieS is build upon a large-scale pretrained transformer backbone (oquab2024dinov2, ; wang2025vggt, ), which encodes each video frame independently and aggregates their information via attentions (vaswani2017attention, ). The aggregated features are then processed by specialized prediction heads: (1) a depth head estimates depth for each input frame, (2) a splatter head predicts per-pixel 3D Gaussian (kerbl20233d, ) appearance attributes, such as color and opacity, for novel view rendering, (3) a motion head estimates the time-conditioned movements of Gaussian primitives towards a target timestamp, allowing us to track its temporal evolution.
Benefiting from the unified architecture, MoVieS can be trained on large-scale datasets featuring both static (zhou2018stereo, ; wang2020tartanair, ; li2023matrixcity, ) and dynamic (mehl2023spring, ; cabon2020virtual, ) scenes, as well as point tracking datasets (zheng2023pointodyssey, ; karaev2023dynamicstereo, ; jin2024stereo4d, ). At inference time, MoVieS takes a monocular video, whether depicting a static or dynamic scene, and reconstructs per-pixel 3D Gaussian primitives along with their motion attributes at any target timestamp, enabling novel view synthesis, depth estimation, and 3D point tracking in a single model.
Extensive experiments on diverse benchmarks (zhou2018stereo, ; yoon2020novel, ; gao2022dynamic, ; koppula2024tapvid, ) show that MoVieS achieves competitive performance across a variety of 4D perception tasks, while being several orders of magnitude faster than the existing state of the art. Furthermore, empowered by novel view synthesis as a proxy task, MoVieS enables dense motion understanding from sparse tracking supervision. This naturally gives rise to a variety of zero-shot applications, further broadening the potential of our approach.
In summary, our main contributions include:
  • We introduce MoVieS, the first feed-forward framework that jointly models appearance, geometry and motion for 4D scene perception from monocular videos.
  • Dynamic splatter pixels are proposed to represent dynamic 3D scenes as renderable deforming 3D particles, bridging novel view synthesis and dynamic geometry reconstruction.
  • MoVieS delivers strong performance and orders of magnitude speedups for 4D reconstruction, and naturally enables a wide range of applications in a zero-shot manner.
Figure 1: Overview. MoVieS consists of a shared image encoder, an attention-based feature backbone (Section 3.2.1), and three heads (Section 3.2.2) that simultaneously predict appearance, geometry and motion. Image shortcut for splatter head and time-varying Gaussian attributes are omitted for brevity.
2 Related Work
2.1 Feed-forward 3D Reconstruction
Most traditional 3D reconstruction methods require dense multi-view supervision and per-scene optimization from scratch at test time (hartley2003multiple, ; davison2007monoslam, ; agarwal2011building, ; mur2015orb, ; schonberger2016structure, ; mildenhall2020nerf, ; kerbl20233d, ). Recent learning-based approaches leverage large-scale data priors to enable feed-forward 3D reconstruction that directly regresses geometric structures such as camera poses and depth (wang2024vggsfm, ; wang2025moge, ; ranftl2020towards, ; piccinelli2024unidepth, ; bochkovskii2024depth, ; yang2024depth, ; hu2024metric3d, ). With the advent of differentiable rendering techniques like 3D Gaussian Splatting (3DGS) (kerbl20233d, ), several works propose to predict pixel-aligned Gaussian attributes for novel view synthesis (charatan2024pixelsplat, ; chen2024mvsplat, ; zhang2024gs, ; ziwen2024long, ; xu2024depthsplat, ). We build on these advances and extend them to dynamic 3DGS representations for novel timestamps.
DUSt3R (wang2024dust3r, ) pioneered direct regression of pixel-aligned pointmaps in a canonical space from image pairs. Subsequent works improved this framework by incorporating feature matching (leroy2024grounding, ), supporting multi-view inputs (wang2025d, ; yang2025fast3r, ; zhang2025flare, ; tang2024mv, ), adapting to streaming video (liu2024slam3r, ; murai2024mast3r, ), or bridging view synthesis with 3DGS (smart2024splatt3r, ; ye2024no, ; zhang2025flare, ; tang2024mv, ). VGGT (wang2025vggt, ) further advances this line by using a stronger image encoder (oquab2024dinov2, ) and multiple task-specific heads to predict all key 3D properties from as few as one to hundreds of views. We adopt VGGT as the backbone of MoVieS for its strong visual representations. However, all these methods are restricted to static scenes and overlook dynamic environments with moving objects—the main focus of our work.
2.2 Dynamic Geometry Reconstruction
Compared to static scene reconstruction, which has been studied for decades, dynamic scene reconstruction remains relatively underexplored. Some works extend the DUSt3R framework to dynamic settings in a plug-and-play manner (chen2025easi3r, ), or leverage foundation models (zhang2024monst3r, ; lu2024align3r, ; yao2025uni4d, ) by incorporating monocular depth (bochkovskii2024depth, ; piccinelli2025unidepthv2, ), optical flow (teed2020raft, ), or 2D point tracking (karaev2024cotracker, ). Dynamic Point Map (sucar2025dynamic, ), a concurrent work, proposes a time-invariant DUSt3R variant by predicting each viewpoint’s pointmap at another timestamp. The most related work, DynaDUSt3R (jin2024stereo4d, ), adds pixel-level motion supervision to model dynamic geometry. However, all these methods are limited to two-frame inputs and produce only point clouds, without high-quality visual reconstructions.
To go beyond two views, CUT3R (wang2025continuous, ) uses a recurrent model to incrementally update scene representations from streaming video for static and dynamic scenes. Other methods (xu2025geometrycrafter, ; jiang2025geo4d, ) treat dynamic reconstruction as conditional video generation, fine-tuning diffusion models to estimate pointmaps probabilistically. However, these require multiple network passes per video.
2.3 Dynamic Novel View Synthesis
Instead of geometric structure, dynamic novel view synthesis emphasizes visual results and view-dependent effects. Early attempts utilize neural radiance field (mildenhall2020nerf, ; fridovich2023k, ; cao2023hexplane, ) to represent dynamic scenes, either from multi-view videos (bansal20204d, ; wang2022fourier, ; li2022neural, ), or monocular ones (pumarola2021d, ; park2021nerfies, ; li2023dynibar, ; zhao2024pgdvs, ). 3DGS (kerbl20233d, ), as a more efficient renderable representation, has been applied for dynamic view synthesis using 4D primitives with a time dimension (yang2024real, ; li2024spacetime, ) or an additional deformable field (wu20244d, ; luiten2024dynamic, ; yang2024deformable, ; lin2024gaussian, ; liu2025modgs, ). Leveraging the low-dimensional nature of scene motion, several works (wang2024shape, ; sun2024splatter, ; stearns2024dynamic, ; lei2025mosca, ; lin2025omniphysgs, ) express motion explicitly and learn Gaussian trajectories across monocular frames. However, these methods are trained from scratch and require iterative optimization for scene adaptation. Off-the-shelf point tracking (doersch2023tapir, ; karaev2024cotracker, ) or optical flow (teed2020raft, ) models are also essential to provide supervision signals.
The most related works to ours are BTimer (liang2024feed, ) and NutWorld (shen2025seeing, ), both of which utilize a feed-forward approach to estimate 3DGS attributes from monocular videos. However, BTimer predicts independent Gaussian chunks per timestamp without modeling frame relations and needs an enhancer module for smooth intermediate frames. NutWorld models Gaussian motion but lacks explicit supervision, relying heavily on pretrained depth (chen2025video, ) and flow (xu2023unifying, ), and uses an orthographic camera, which may further lead to projection distortions.
3 Method
Following previous studies on monocular 4D reconstruction (zhao2024pgdvs, ; wang2024shape, ; liu2025modgs, ; lei2025mosca, ; liang2024feed, ), given a posed video $\mathcal{V}\coloneqq\{\mathbf{I}_{i},\mathbf{P}_{i},\mathbf{K}_{i}\}_{i=1}^{N}$ with frames $\mathbf{I}_{i}\in\mathbb{R}^{3\times H\times W}$, camera poses $\mathbf{P}_{i}\in\mathbb{SE}(3)$, intrinsics $\mathbf{K}_{i}\in\mathbb{R}^{3\times 3}$ and timestamps $t_{i}\in[0,1]$, we aim to train a generic model that jointly reconstruct the underlying appearance, geometry and motion of dynamic scenes.
3.1 Dynamic Splatter Pixel
To model 3D scenes with moving contents, we propose a novel representation, namely dynamic splatter pixel, which decomposes dynamic scenes into a set of static Gaussian primitives and their corresponding deformation fields. Given an input video $\mathcal{V}$, each pixel in the $i$-th frame $\mathbf{I}_{i}$ is associated with a splatter pixel $\mathbf{g}_{i}$ (charatan2024pixelsplat, ; szymanowicz2024splatter, ; zhang2024gs, ) in a shared canonical space of the first frame camera’s coordinate system. Each $\mathbf{g}\coloneqq\{\mathbf{x},\mathbf{a}\}$ is parameterized by its position $\mathbf{x}\in\mathbb{R}^{3}$ in the canonical space and other rendering attributes $\mathbf{a}\in\mathbb{R}^{11}$, including rotation quaternion $\mathbf{q}\in\mathbb{R}^{4}$, scale $\mathbf{s}\in\mathbb{R}^{3}$, opacity $\alpha\in\mathbb{R}$, and color $\mathbf{c}_{\text{rgb}}\in\mathbb{R}^{3}$ (kerbl20233d, ). Considering splatter pixels were originally designed for static scenes, we decouple motion from geometric structure to adapt them for dynamic scenes. An additional time-dependent deformation field is introduced, in which each splatter pixel $\mathbf{g}$ is associated with $\mathbf{m}(t)\coloneqq\{\Delta\mathbf{x}(t),\Delta\mathbf{a}(t)\}$. $\Delta\mathbf{x}(t)\in\mathbb{R}^{3}$ is the motion vector of a splatter pixel at time $t$ with respect to the canonical 3D space, and $\Delta\mathbf{a}(t)$ is the change of the corresponding attributes of the splatter pixel at time $t$. Therefore, the splatter pixel $\mathbf{g}$ is deformed to time $t$ as:
$\mathbf{x}\leftarrow\mathbf{x}+\Delta\mathbf{x}(t),\quad\mathbf{a}\leftarrow\mathbf{a}+\Delta\mathbf{a}(t).$ (1)
By combining static splatter pixels and their deformation fields, we establish the correspondence between each Gaussian primitive and its temporal dynamics, thereby enabling dynamic scene modeling and dense motion estimation. In practice, we find that making only the scale and rotation attributes time-dependent, i.e., $\Delta\mathbf{a}(t)\in\mathbb{R}^{7}$, is sufficient for representing dynamic splatter pixels.
Figure 2: Motion Head. Given $t_{q}$ target timesteps, the proposed motion head is conditioned via adaptive layer normalization (AdaLN) and predicts 3D displacements for each input pixel. After rasterization using the $M$ corresponding query-time cameras, output images in shape $M\times 3\times H\times W$ are rendered for supervision. Gaussian attribute deformation $\Delta\mathbf{a}$ is omitted for brevity.
3.2 MoVieS: Unify Appearance, Geometry and Motion
As illustrated in Figure 1, the framework of the proposed MoVieS provides a unified approach to simultaneously model appearance, geometry and motion of dynamic scenes. It consists of a feature backbone with camera and time information for extracting features from input video frames, followed by dedicated heads for depth, splatter and motion estimation.
3.2.1 Feature Backbone
Given an input posed video $\mathcal{V}=\{\mathbf{I}_{i},\mathbf{P}_{i},\mathbf{K}_{i},t_{i}\}_{i=1}^{N}$, we first patchify the input images $\mathbf{I}_{i}$ and use a pretrained image encoder (oquab2024dinov2, ) to extract their features. To effectively incorporate camera information, we adopt two complementary strategies to embed the camera parameters into image features: (1) Plücker embedding: camera pose $\mathbf{P}_{i}$ and intrinsics $\mathbf{K}_{i}$ are encoded into pixel-aligned Plücker embeddings (sitzmann2021light, ), which are then downsampled and fused with image features via spatial-wise addition; (2) Camera token: both $\mathbf{P}_{i}$ and $\mathbf{K}_{i}$ are passed through a linear layer to generate a camera token, which is appended to the sequence of image tokens. Plücker embedding provides a dense and spatially aligned encoding of camera geometry, while the camera token interacts globally with image tokens through attention, enabling more holistic feature reasoning. Ablation study on these two camera injection manners is validated in Section 4.4.
To inform the model that input images originate from a temporally ordered video, we additionally encode each timestamp $t_{i}\in[0,1]$ using sinusoidal positional encoding (mildenhall2020nerf, ) to produce a time token, which is then concatenated with the aforementioned image and camera tokens.
After tokenizing input images, camera parameters, and timestamps, we apply geometrically pretrained attention blocks (vaswani2017attention, ) from VGGT (wang2025vggt, ) to enable interactions among image tokens across video frames. This produces a set of shared feature tokens enriched with inter-frame context as well as camera and temporal information, which are then used for predicting various properties of dynamic scenes.
3.2.2 Prediction Heads
Shared aggregated video tokens from the feature backbone are fed into three parallel prediction heads that estimate the appearance, geometry, and motion of the dynamic scene, respectively. Each head adopts a DPT-style architecture (ranftl2021vision, ) to convert image tokens into dense predictions matching the input resolution, thereby producing dynamic splatter pixels.
Depth and Splatter Head
Different from previous feed-forward 3DGS reconstruction methods (charatan2024pixelsplat, ; chen2024mvsplat, ; zhang2024gs, ) that use a single head to predict all splatter pixel attributes, we adopt a decoupled design to better leverage geometric priors from the pretrained VGGT (wang2025vggt, ). Specifically, a dedicated depth head, initialized from VGGT, is used for geometry prediction to provide spatial grounding for splatter pixel construction, while a separate DPT as splatter head is trained from scratch for appearance rendering. We further incorporate a direct RGB shortcut (ye2024no, ) from the input image to the final convolution layer of the splatter head to preserve high-frequency details and enhance color fidelity.
Motion Head
To capture scene dynamics, a novel motion head as shown in Figure 2 is introduced to predict dense deformation $\mathbf{m}(t)$ for each dynamic splatter pixel at any target moment. Temporal variation is enabled by injecting the sinusoidally encoded query time $t_{q}$ into the image tokens via adaptive layer normalization (xu2019understanding, ) before applying DPT convolutions. For each input frame $\mathbf{I}_{i}$ at time $t_{i}$, the motion head predicts its 3D movements toward $t_{q}$ in a shared world coordinate system. Thus, for $N$ input frames and $M$ query timestamps, this yields a motion map of shape $M\times N\times 3\times H\times W$.
3.3 Training
Table 1: Training Datasets. Nine datasets from diverse sources are utilized to train MoVieS at scale. “#Repeat” denotes dataset duplication count during integration to balance their contributions.
Dataset Dynamic? Depth? Tracking? Real? #Scenes #Frames #Repeat
RealEstate10K (zhou2018stereo, ) 70K 6.36M 1$\times$
TartanAir (wang2020tartanair, ) 0.4K 0.49M 100$\times$
MatrixCity (li2023matrixcity, ) 4.5K 0.31M 10$\times$
PointOdyssey (zheng2023pointodyssey, ) 0.1K 0.18M 1000$\times$
DynamicReplica karaev2023dynamicstereo 0.5K 0.26M 100$\times$
Spring (mehl2023spring, ) 0.03K 0.003M 2000$\times$
VKITTI2 (cabon2020virtual, ) 0.1K 0.03M 500$\times$
Stereo4D (jin2024stereo4d, ) 98K 19.6M 1$\times$
3.3.1 Dataset Curation
An ideal dataset for dynamic scene reconstruction would include synchronized multi-view videos with dense depth and point tracking annotations. However, such data is infeasible to capture and annotate at scale in practice. Instead, we leverage a diverse set of open-source datasets (zhou2018stereo, ; wang2020tartanair, ; li2023matrixcity, ; mehl2023spring, ; cabon2020virtual, ; zheng2023pointodyssey, ; karaev2023dynamicstereo, ; jin2024stereo4d, ), each providing complementary supervision, as shown in Table 1. With the flexible model design, MoVieS can be jointly trained on these heterogeneous sources by aligning objectives to their respective annotations. More details about data curation are provided in Appendix A.
3.3.2 Objectives
MoVieS is trained by a multi-task objective that combines depth, rendering and motion losses:
$\mathcal{L}\coloneqq\lambda_{\text{d}}\mathcal{L}_{\text{depth}}+\lambda_{\text{r}}\mathcal{L}_{\text{rendering}}+\lambda_{\text{m}}\mathcal{L}_{\text{motion}}.$ (2)
Depth and Rendering Losses
Depth loss is computed as the mean squared error (MSE) between the predicted and ground truth depth maps, along with their spatial gradients, after filtering out invalid values. Rendering loss combines pixel-wise MSE and perceptual loss (zhang2018perceptual, ) between 3DGS-rendered images under corresponding camera views and the video frames at target timestamps.
Motion Loss
Given 3D point tracking datasets, ground-truth motion $\Delta\mathbf{x}$ is defined as the 3D displacement of each tracked point between any two frames. Since all 3D points are defined in the world coordinate and most tracked points remain static, implying that their corresponding motion vectors tend to zero. We apply a point-wise L1 loss between the predicted and ground-truth motions to promote sparsity after filtering out points that are not visible in the input frames. Additionally, to complement direct point-to-point alignment, a distribution loss is introduced that encourages the predicted motion vectors to preserve the internal relative distance structure within each frame. The final motion loss is defined as a combination of point-wise and distribution-level supervision:
$\mathcal{L}_{\text{motion}}$ $\coloneqq\lambda_{\text{pt}}\mathcal{L}_{\text{pt}}+\lambda_{\text{dist}}\mathcal{L}_{\text{dist}}$ (3)
$\ =\frac{1}{P}\sum_{i\in\Omega}\lambda_{\text{pt}}\|\Delta\hat{\mathbf{x}}_{i}-\Delta\mathbf{x}_{i}\|_{1}+\frac{1}{P^{2}}\sum_{(i,j)\in\Omega\times\Omega}\lambda_{\text{dist}}\|\Delta\hat{\mathbf{x}}_{i}\cdot\Delta\hat{\mathbf{x}}_{j}^{\top}-\Delta\mathbf{x}_{i}\cdot\Delta\mathbf{x}_{j}^{\top}\|_{1},$ (4)
where $\Omega$ denotes the set of all valid $P$ tracked points, and $\Omega\times\Omega$ is its Cartesian product. Ablation study on the effectiveness of these two types of motion supervision is presented in Section 4.4.
Normalization
Similar to VGGT (wang2025vggt, ), we normalize the 3D scene scale by the average Euclidean distance from each 3D point to the origin of the canonical world coordinate system. As a result, unlike some other reconstruction methods (wang2024dust3r, ; leroy2024grounding, ; wang2025continuous, ), we do not apply additional normalization in the depth or motion loss. We also omit confidence-aware weighting (wang2024dust3r, ; wang2025vggt, ) for simplicity.
4 Experiments
4.1 Experimental Settings
Implementation
MoVieS is built on a geometrically pretrained transformer, VGGT (wang2025vggt, ), with splatter head and camera/time embeddings trained from scratch. AdamW (loshchilov2017decoupled, ) with cosine learning rate scheduling and linear warm-up is used for optimization. We observed that training MoVieS is particularly unstable, likely arising from sparse annotations and the heterogeneous nature of the training data. A curriculum strategy is employed that gradually increases the training complexity, involving (1) pretraining on static scenes, (2) dynamic scenes with varying views, and (3) fine-tuning on high resolution. Several techniques, such as gsplat (ye2025gsplat, ) rendering backend, DeepSpeed (rasley2020deepspeed, ), gradient checkpointing (chen2016training, ), gradient accumulation, and bf16 mixed precision, are adopted to improve memory and computation efficiency. Training completes in approximately 5 days using 32 H20 GPUs. Please refer to Appendix A for more details about implementations.
Evaluation
We evaluate the capabilities of MoVieS on two primary tasks: novel view synthesis (Section 4.2) and 3D point tracking (Section 4.3). Following prior works (charatan2024pixelsplat, ; chen2024mvsplat, ; xu2024depthsplat, ; wang2024shape, ; sun2024splatter, ; lei2025mosca, ; liang2024feed, ), RealEstate10K (zhou2018stereo, ) is used to evaluate novel view synthesis performance on static scenes. DyCheck (gao2022dynamic, ) and NVIDIA dynamic scene dataset (yoon2020novel, ) are adopted for dynamic scenes, in terms of PSNR, SSIM and LPIPS (zhang2018perceptual, ). Since the DyCheck dataset contains invisible regions in the target novel views, we compute reconstruction metrics (denoted by a prefix “m”) using the provided covisibility masks, ensuring a fair comparison across methods. For 3D point tracking, the TAPVid-3D (koppula2024tapvid, ) benchmark serves as the evaluation protocol, covering both indoor and outdoor scenes across three datasets. End-point Euclidean error in the 3D space (EPE3D) and the percentage of 3D points within $0.05$ and $0.10$ units of the ground-truth locations ($\delta^{0.05}_{3D}$ and $\delta^{0.10}_{3D}$) are utilized as metrics.
4.2 Novel View Synthesis
4.2.1 Static Scene View Synthesis
As a special case of dynamic scenes, we first evaluate MoVieS on a static dataset RealEstate10K (zhou2018stereo, ), comparing against several state-of-the-art feed-forward static scene reconstruction methods, including DepthSplat (xu2024depthsplat, ), our reimplementation of GS-LRM (zhang2024gs, ), and our method pretrained solely on the first-stage static reconstruction. As shown in Table 2, although MoVieS is primarily designed for dynamic scenes, it maintains competitive performance on static scenes. Notably, when processing static inputs, our predicted motion naturally converges to zero, demonstrating the ability of MoVieS to implicitly differentiate between static and dynamic regions without explicit supervision.
4.2.2 Dynamic Scene View Synthesis
Dynamic scene view synthesis is evaluated on two benchmarks: (1) DyCheck (gao2022dynamic, ), with 7 daily videos captured by 3 synchronized cameras, and (2) NVIDIA (yoon2020novel, ), containing 9 dynamic scenes recorded by a 12-camera static rig. For DyCheck, we sample 13 frames of size $518\times 518$ from a random clip from iPhone-record videos, and include the other two cameras for evaluation. For NVIDIA, 12 input views at $379\times 672$ resolution are sampled in a round-robin style (gao2021dynamic, ; liu2023robust, ; liang2024feed, ), where the i-th frame comes from the i-th camera at timestep i. We compare against three state-of-the-art optimization-based methods (wang2024shape, ; sun2024splatter, ; lei2025mosca, ) for dynamic 3D Gaussian reconstruction, and also report static feed-forward baselines for reference.
As shown in Table 2, MoVieS achieves competitive or superior performance compared to baselines, while requiring only 0.93s per scene, which is orders of magnitude faster than prior approaches that rely on heavy pretrained models and complex multi-stage pipelines. Qualitative visualizations in Figure 3 further highlight the strength of our approach. While MoSca (lei2025mosca, ) demonstrates impressive performance, it struggles with sparse inputs and often overfits to seen poses, producing spiky and over-saturated artifacts under novel views and timesteps. In contrast, MoVieS leverages large-scale learned priors to generalize more effectively, yielding smoother and more realistic results.
To ensure fair comparison and better reflect real-world scenarios, no video masks for dynamic objects are used in our experiments. It poses a major challenge for optimization-based methods like Shape-of-Motion (wang2024shape, ) and Splatter-a-Video (sun2024splatter, ), which rely heavily on explicit motion segmentation. The challenge is particularly evident on the NVIDIA dataset, where significant camera shake hinders disentangling scene dynamics, leading to notably degraded performance and even worse than static baselines. In contrast, our method exhibits strong robustness by directly learning to model motion.
Figure 3: Novel View Synthesis for Dynamic Scenes. Given a monocular video, we compare synthesized views from a novel view across different methods. Regions invisible in the input are rendered as black or white, depending on the rendering implementation. More results in Figure 6.
Table 2: Evaluation on Novel View Synthesis. The best, second best and third best results are highlighted for clarity. $\dagger$ indicates our reimplemented version of GS-LRM (zhang2024gs, ).
Novel View Synthesis Time Per Scene RealEstate10K DyCheck NVIDIA
$\uparrow$ PSNR $\uparrow$ SSIM $\downarrow$ LPIPS $\uparrow$ mPSNR $\uparrow$ mSSIM $\downarrow$ mLPIPS $\uparrow$ PSNR $\uparrow$ SSIM $\downarrow$ LPIPS
Static Feed-forward
DepthSplat (xu2024depthsplat, ) 0.60s 26.57 80.06 0.124 13.83 0.4364 0.3850 17.16 0.5002 0.3023
GS-LRM (zhang2024gs, ) 0.57s 26.94 79.13 0.139 14.60 0.4535 0.3775 17.83 0.4988 0.3265
Ours (static) 0.84s 27.60 81.25 0.113 15.24 0.4784 0.3783 18.73 0.5042 0.2959
Optimization-based
Splatter-a-Video (sun2024splatter, ) 37min - - - 13.61 0.3131 0.5706 14.39 0.2538 0.5983
Shape-of-Motion (wang2024shape, ) 10min - - - 17.96 0.5662 0.3463 15.30 0.3169 0.5087
MoSca (lei2025mosca, ) 45min - - - 18.24 0.5514 0.3698 21.45 0.7123 0.2653
Ours 0.93s 26.98 81.75 0.111 18.46 0.5887 0.3094 19.16 0.5141 0.3152
4.3 3D Point Tracking
Trained on large-scale point tracking datasets (zheng2023pointodyssey, ; karaev2023dynamicstereo, ; jin2024stereo4d, ), the proposed method can also densely track any 3D point corresponding to a pixel across video frames (see Figure 7 in the appendix). We compare MoVieS against three strong baselines, including two state-of-the-art 2D point tracking methods, BootsTAP (doersch2024bootstap, ) and CoTracker3 (karaev2024cotracker, ), as well as a native 3D point tracking approach, SpatialTracker (xiao2024spatialtracker, ). For 2D trackers, a recent video depth estimation model (chen2025video, ) and ground-truth camera intrinsics are used to unproject tracked points into 3D space. To account for scale differences across methods, we normalize all predicted 3D points by their median norm before evaluation.
Quantitative results are reported in Table 3. While 3D-based SpatialTracker generally outperforms 2D-based approaches, all of them rely heavily on pretrained monocular depth estimators for geometry reasoning, introducing significant noise and inconsistency in the 3D space. In contrast, MoVieS directly estimates 3D point positions in a shared world coordinate, enabling more accurate and robust 3D tracking, and achieves consistently superior or competitive performance across all datasets.
Table 3: Evaluation on 3D Point Tracking. The best, second best and third best results are highlighted for clarity. $\dagger$ denotes 3D points obtained by combining a depth estimation model (chen2025video, ).
3D Point Tracking Aria Digital Twin DriveTrack Panoptic Studio
$\downarrow$ EPE3D $\uparrow$ $\delta^{0.05}_{3D}$uparrow $\delta^{0.10}_{3D}$downarrow EPE3D $\uparrow$ $\delta^{0.05}_{3D}$uparrow $\delta^{0.10}_{3D}$downarrow EPE3D $\uparrow$ $\delta^{0.05}_{3D}$uparrow $\delta^{0.10}_{3D}$
BootsTAPIR (doersch2024bootstap, ) 0.5539 17.73% 32.97% 0.0617 55.82% 75.66% 0.0650 69.28% 87.95%
CoTracker3 (karaev2024cotracker, ) 0.5614 19.88% 35.82% 0.0637 55.30% 77.55% 0.0617 69.27% 88.04%
SpatialTracker (xiao2024spatialtracker, ) 0.5413 18.08% 38.23% 0.0648 56.58% 80.67% 0.0519 72.91% 89.86%
Ours 0.2153 52.05% 71.63% 0.0472 60.63% 79.87% 0.0352 87.88% 94.61%
4.4 Ablation and Analysis
Figure 4: Motion Visualization for Ablation Studies. We investigate key factors affecting motion learning in MoVieS, such as loss design and synergy with view synthesis. XYZ values in motion maps are normalized as RGB for visualization. Red arrows on video frames indicate motion directions.
Table 4: Ablation study on camera conditioning.
Camera Conditioning RealEstate10K
$\uparrow$ PSNR $\uparrow$ SSIM $\downarrow$ LPIPS
N/A 25.56 74.13 0.150
+ Plücker embedding 25.81 74.44 0.143
+ Camera token 26.81 78.65 0.121
Ours (static) 27.60 81.25 0.113
Table 5: Ablation study on motion supervision.
Motion Supervision Aria Digital Twin
$\downarrow$ EPE3D $\uparrow$ $\delta^{0.05}_{3D}$uparrow $\delta^{0.10}_{3D}$
N/A 0.7938 19.58% 32.86%
+ Point-wise L1 0.2262 48.74% 69.93%
+ Distribution loss 0.2496 45.98% 66.87%
Ours 0.2153 52.05% 71.63%
Table 6: Ablation Study on the synergy of motion estimation and novel view synthesis (NVS).
Motion & View Synthesis DyCheck NVIDIA Aria Digital Twin
$\uparrow$ mPSNR $\uparrow$ mSSIM $\downarrow$ mLPIPS $\uparrow$ PSNR $\uparrow$ SSIM $\downarrow$ LPIPS $\downarrow$ EPE3D $\uparrow$ $\delta^{0.05}_{3D}$uparrow $\delta^{0.10}_{3D}$
NVS w/o motion 15.82 0.4796 0.3741 18.38 0.4856 0.3010 0.7938 19.58% 32.86%
Motion w/o NVS 16.26 0.4556 0.3461 18.98 0.4939 0.3207 0.3801 24.72% 42.92%
Ours 18.46 0.5887 0.3094 19.16 0.5041 0.3152 0.2153 52.05% 71.63%
Camera conditioning
We investigate different camera conditioning strategies in the static pertaining stage. Quantitative comparisons are provided in Table 5. Camera tokens are injected throughout the feature backbone, enabling effective camera-aware modeling. In contrast, Plücker embeddings provide limited conditioning on their own and are merely comparable to having no camera information at all. However, as a pixel-aligned representation, Plücker embeddings are complementary to camera tokens, and their combination yields the most effective camera conditioning.
Motion Supervision
To learn object 3D movements in dynamic scenes, we provide two kinds of motion supervision: (1) point-wise L1 loss and (2) distribution loss, as described in Equation 4. Their efficiency is evaluated via the 3D point tracking task in Table 5. Without any motion supervision, i.e., learning solely from novel view synthesis, training exhibits severe loss oscillations and frequent None gradients. The distribution loss captures only relative motion between pixels, while point-wise L1 loss produces more reasonable motion maps. Combining both leads to sharper boundaries. Qualitative results of estimated motions from different motion objectives are provided in Figure 4.
Synergy of Motion and View Synthesis
Thanks to the unified design of MoVieS, it supports simultaneous novel view synthesis (NVS) and motion estimation. We study their mutual benefits in Table 6. “NVS w/o motion” disables explicit motion supervision during training, relying solely on NVS as a proxy to learn dynamics. As shown in Table 6 and Figure 4, this setting fails to learn meaningful motion and tends to model static scenes. “Motion w/o NVS” detaches the motion head from 3DGS rendering and instead conditions the depth head on time. Although explicit supervision enables some motion learning, the predictions are blurry and low-quality, as shown in Figure 4. Moreover, the depth head must now model both geometry and dynamics, increasing its burden and negatively affecting NVS. These results highlight the mutual reinforcement between NVS and motion estimation in MoVieS, where joint training leads to better performance on both.
4.5 Zero-shot Applications
Scene Flow Estimation
Scene flow can be naturally derived by transforming the estimated per-pixel motion vectors from world coordinates to the target camera’s coordinates. Visualization results in Figure 5(a) present sharp edges and accurate motion directions. More results are provided in Figure 8.
Moving Object Segmentation
By thresholding the norm of per-pixel motion vectors, the estimated motion maps can be used to segment moving objects (Figure 5(b)), which is an essential task in computer vision and robotics (xie2024moving, ; huang2025segment, ). Remarkably, this is achieved without any explicit mask supervision, demonstrating the strong potential of our method. More results are provided in Figure 9.
Figure 5: Zero-shot Applications. The predicted pixel-aligned motion maps from our model can be directly applied to downstream tasks, such as (a) scene flow estimation and (b) moving object segmentation, in a zero-shot manner, without any task-specific fine-tuning or supervision.
5 Conclusion
In this work, we introduce MoVieS, a feed-forward model for dynamic novel view synthesis from monocular videos. It’s trained on large-scale datasets from diverse sources and jointly models scene appearance, geometry and motion in a unified and efficient network. Dynamic splatter pixels are proposed to represent dynamic scenes, enabling accurate and temporally coherent 4D reconstruction. Beyond novel view synthesis, MoVieS supports a broad range of applications including depth estimation, 3D point tracking, scene flow estimation and moving object segmentation, demonstrating its versatility for dynamic scene perception. We hope this work could serve as a step toward generalizable dynamic scene understanding, and support applications that require spatial and motion intelligence. More discussions on limitations, future work and broader impact are included in Appendix B.
Appendix A Implementation Details
A.1 Dataset Details
As summarized in Table 1, the majority of training datasets used in our experiments are synthetic, providing rich annotations such as pixel-aligned depth maps and accurate camera intrinsics and extrinsics. Ground-truth 3D point trajectories are also available in PointOdyssey [33] and DynamicReplica [34]. Stereo4D [35], on the other hand, is a large-scale dataset constructed from YouTube stereo videos and annotated using pretrained foundation models [62, 98]. Despite being partially noisy, its diversity and scale offer strong generalization benefits.
The utilized backbone VGGT [10] requires all 3D scenes to be normalized to a unit scale, which in turn necessitates depth maps and camera extrinsics to be defined in a consistent metric space. This requirement is satisfied by the synthetic datasets and Stereo4D, whose camera poses and depths are metric-aligned by construction. In contrast, RealEstate10K [28] only provides relative camera parameters estimated via COLMAP [6], resulting in an unknown global scale. To address this, we re-estimate both depth maps and camera extrinsics using recent foundation models, Video Depth Anything [5] and MegaSaM [101], to recover aligned geometry across frames.
A.2 Curriculum Training
We observed that training MoVieS is particularly unstable: the training loss often fluctuates abruptly, and gradients are prone to becoming None. It may arise from sparse annotations and the heterogeneous nature of the training data, which mixes datasets from various sources with differing domains (e.g., indoor vs. outdoor, real vs. synthetic), camera fields of view, recording frame rates, etc.
Thanks to the versatile design of MoVieS, we employ a curriculum strategy that gradually increases the training complexity. It begins by pretraining the model on low-resolution ($224\times 224$) static datasets with only depth and photometric losses, then introduces dynamic datasets along with motion supervision. We found that static datasets play a crucial role in stabilizing training for dynamic scenes, without which the loss would be highly unstable. Since modeling dynamic scenes requires reconstructing a set of 3DGS for each query time, which results in high GPU memory usage, we start by training on 5 input views for dynamic scenes and then expand to 13 views for fine-tuning. Finally, the training resolution is increased to $518\times 518$.
A.3 Training Details
Image encoder, feature backbone and depth head of MoVieS are initialized from a geometrically pretrained transformer, VGGT [10]. Motion head is initialized from its point head. Remaining components, such as the splatter head and camera/time embeddings, are trained from scratch. We use AdamW optimizer [92] with a weight decay of 0.05, and adopt a cosine learning rate scheduler [102] with linear warm-up for all curriculum training stages. For static pretraining and dynamic scenes with 5 and 13 input views at a resolution of $224\times 224$, we use learning rates of 4e-4, 4e-4 and 4e-5, respectively, with a batch size of 256. Training rates for parameters initialized from VGGT are multiplied by 0.1. With 32 H20 GPUs, training takes about 2 days for static and then dynamic scenes with 5 views, and 2 days for 13 views. MoVieS is then finetuned on $518\times 518$ videos with 13 frames using a learning rate of 1e-5, which takes around 1 day. To improve memory and computation efficiency, several techniques such as gsplat [93] rendering backend, DeepSpeed [94], gradient checkpointing [95], gradient accumulation, and bf16 mixed precision are adopted in this work.
Appendix B Discussion
B.1 Limitations and Future Work
Although MoVieS achieves competitive performance with inference speeds that are orders of magnitude faster than optimization-based methods, there remains a noticeable gap in reconstruction quality. This gap arises partly because many optimization-based approaches benefit from multiple pretrained models that preprocess input videos to provide richer and more accurate cues. Incorporating such richer prior knowledge directly into MoVieS represents a promising direction for future work to further improve reconstruction fidelity and robustness.
Currently, MoVieS depends on off-the-shelf tools for camera parameter estimation, which adds an external dependency and may limit end-to-end optimization. Seamlessly integrating camera pose estimation within the MoVieS pipeline could simplify the overall workflow, reduce error accumulation, and enhance adaptability to diverse scenarios.
Moreover, scaling MoVieS to handle large-scale scenes and achieve high-resolution rendering remains a challenge. The computational cost and memory demand grow significantly with scene complexity and resolution, which constrains practical deployment. Developing more compact and efficient dynamic scene representations, possibly through novel model architectures or sparse encoding strategies, is essential to push 4D reconstruction towards real-world applications.
Addressing these challenges will not only bridge the performance gap but also unlock the full potential of fast and high-quality dynamic scene reconstruction.
B.2 Broader Impact
MoVieS provides substantial advantages for fields including robotics simulation, AR/VR, autonomous driving, and digital twins by enabling fast and generalizable dynamic scene understanding. Its ability to reconstruct dynamic environments efficiently can accelerate innovation and improve system performance in these areas. However, such powerful technology also raises risks related to the unauthorized generation of content and potential privacy violations. To mitigate these risks, it is essential to establish clear ethical guidelines and implement appropriate regulatory measures to ensure the responsible and safe use of this technology.
Appendix C More Visualization Results
We provide more visualization results of MoVieS in Figure 6, 7, 8 and 9.
Appendix D License Information
We employ several open-source implementations in our experimental comparisons, including: (1) DepthSplat [24] [^1] https://github.com/cvg/depthsplat/tree/main (MIT License), (2) Splatter-a-Video [81] [^2] https://github.com/SunYangtian/Splatter_A_Video (Apache License), (3) Shape-of-Motion [80] [^3] https://github.com/vye16/shape-of-motion/ (MIT License), (4) MoSca [83] [^4] https://github.com/JiahuiLei/MoSca (MIT License), (5) BootsTAPIR [98] [^5] https://github.com/google-deepmind/tapnet (Apache License), (6) CoTracker3 [19] [^6] https://github.com/facebookresearch/co-tracker (Creative Commons Attribution-NonCommercial 4.0 International Public License), and (7) SpatialTracker [76] [^7] https://github.com/henry123-boy/SpaTracker (Attribution-NonCommercial 4.0 International).
Several public datasets from diverse sources are utilized to train our models, including: (1) RealEstate10K [28] [^8] https://google.github.io/realestate10k/ (Creative Commons Attribution 4.0 International License), (2) TartanAir [29] [^9] https://theairlab.org/tartanair-dataset/ (Creative Commons Attribution 4.0 International License), (3) MatrixCity [30] [^10] https://city-super.github.io/matrixcity/ (Apache License), (4) PointOdyssey [33] [^11] https://pointodyssey.com/ (MIT License), (5) DynamicReplica [34] [^12] https://github.com/facebookresearch/dynamic_stereo (Attribution-NonCommercial 4.0 International License), (6) Spring [31] [^13] https://spring-benchmark.org/ (CC BY 4.0), (7) VKITTI2 [32] [^14] https://europe.naverlabs.com/research/computer-vision/proxy-virtual-worlds-vkitti-2/ (Creative Commons Attribution-NonCommercial-ShareAlike 3.0), and (8) Stereo4D [35] [^15] https://stereo4d.github.io/ (CC0 1.0 Universal).
Figure 6: Qualitative results for reconstruction. “Motion” means the 3D points’ movements of the input frame with respect to the first one.
Figure 7: Qualitative results for 3D point tracking. “Motion” means the 3D points’ movements of the input frame with respect to the first one.
Figure 8: More results for scene flow estimation. “Flow” means the optical flow of the input pixels with respect to the first frame. Arrows on pictures roughly mark directions.
Figure 9: More results for moving object segmentation. Masks for moving parts, which are filtered from motion maps, are highlighted across frames in light red.

### v2

@@
 Abstract
-We present MoVieS, a novel feed-forward model that synthesizes 4D dynamic novel views from monocular videos in one second.
-MoVieS represents dynamic 3D scenes using pixel-aligned grids of Gaussian primitives, explicitly supervising their time-varying motion.
-This allows, for the first time, the unified modeling of appearance, geometry and motion, and enables view synthesis, reconstruction and 3D point tracking within a single learning-based framework.
-By bridging novel view synthesis with dynamic geometry reconstruction, MoVieS enables large-scale training on diverse datasets with minimal dependence on task-specific supervision.
+We present MoVieS, a Motion-aware View Synthesis model that reconstructs 4D dynamic scenes from monocular videos in one second.
+It represents dynamic 3D scenes with pixel-aligned Gaussian primitives and explicitly supervises their time-varying motions.
+This allows, for the first time, the unified modeling of appearance, geometry and motion from monocular videos, and enables reconstruction, view synthesis and 3D point tracking within a single learning-based framework.
+By bridging view synthesis with geometry reconstruction, MoVieS enables large-scale training on diverse datasets with minimal dependence on task-specific supervision.
 As a result, it also naturally supports a wide range of zero-shot applications, such as scene flow estimation and moving object segmentation.
 Extensive experiments validate the effectiveness and efficiency of MoVieS across multiple tasks, achieving competitive performance while offering several orders of magnitude speedups.
-[…]
+{NoHyper}: Equal contribution; $\dagger$: Project lead; $\ddagger$: Corresponding author.
 1 Introduction
-Humans and animals perceive a continuous stream of observations from a dynamic three-dimensional world, and effortlessly interpret its underlying 3D geometry and motion.
+Humans and animals perceive a continuous stream of observations from a dynamic 3D world, and effortlessly interpret its underlying geometry and motion.
 Replicating this capability is essential for any embodied agent that must understand and act in the physical world.
-Recent advances have made great strides in individual 3D tasks, such as monocular depth estimation (ranftl2020towards, ; ke2024repurposing, ; yang2024depth, ; bochkovskii2024depth, ; chen2025video, ), 3D scene reconstruction (schonberger2016structure, ; teed2021droid, ; wang2024dust3r, ; zhang2024monst3r, ; wang2025vggt, ), novel view synthesis (mildenhall2020nerf, ; fridovich2023k, ; kerbl20233d, ; wu20244d, ; luiten2024dynamic, ) and point tracking (harley2022particle, ; doersch2023tapir, ; wang2023tracking, ; karaev2024cotracker, ; xiao2024spatialtracker, ).
+Recent advances have made great strides in individual 3D tasks, such as monocular depth estimation [54, 23, 84, 2, 6], 3D scene reconstruction [56, 65, 74, 91, 68], novel view synthesis [45, 12, 24, 76, 43] and point tracking [15, 11, 71, 22, 77].
 However, besides treating each task in isolation, most existing view synthesis and reconstruction studies focus on static scenes and require costly per-scene optimization without learning prior knowledge.
 Real-world environments are inherently dynamic and diverse, and all the aforementioned 3D scene understanding tasks could share common underlying principles.
-Motivated by this, we introduce MoVieS, a Motion-aware dynamic View Synthesis model for feed-forward 4D reconstruction of videos in the wild, that jointly models scenes’ appearance, geometry and motion.
-MoVieS represents 3D dynamic scenes using renderable and deformable 3D particles, termed dynamic splatter pixels, and utilizes a differentiable 3D Gaussian rendering framework (kerbl20233d, ).
-Specifically, following recent practices on feed-forward view synthesis (charatan2024pixelsplat, ; szymanowicz2024splatter, ; zhang2024gs, ; xu2024depthsplat, ; lin2025diffsplat, ), each input pixel is mapped to a 3D Gaussian primitive, with its 3D location determined by predicted depth.
+Figure 1: Overview.
+MoVieS consists of a shared image encoder, an attention-based feature backbone (Sec.
+3.2.1), and three heads (Sec.
+3.2.2) to jointly model appearance, geometry and motion.
+Motion head is time-conditioned to model dynamic content with respect to several query timestamps.
+Normalized XYZ values in the 3D space of motion maps are treated as RGB channels for visualization.
+Time-varying Gaussian attributes are omitted and point clouds with color and identity Gaussian attributes are visualized here for brevity.
+Motivated by this, we introduce MoVieS, a Motion-aware dynamic View Synthesis model for feed-forward 4D reconstruction of monocular videos, that jointly models scenes’ appearance, geometry and motion.
+MoVieS represents 3D dynamic scenes using renderable and deformable 3D particles, termed dynamic splatter pixels, and utilizes a differentiable 3D Gaussian rendering framework [24].
+Specifically, following recent practices on feed-forward view synthesis [5, 63, 92, 79, 34], each input pixel is mapped to a 3D Gaussian primitive, with its 3D location determined by predicted depth.
 To model dynamics, MoVieS regresses per-pixel motion displacements toward arbitrary query timestamps, enabling temporal tracking of each splatter pixel.
 This design facilitates coherent reconstruction of both 3D geometry and appearance across camera viewpoints and temporal frames.
-As shown in Figure 1, MoVieS is build upon a large-scale pretrained transformer backbone (oquab2024dinov2, ; wang2025vggt, ), which encodes each video frame independently and aggregates their information via attentions (vaswani2017attention, ).
-The aggregated features are then processed by specialized prediction heads: (1) a depth head estimates depth for each input frame, (2) a splatter head predicts per-pixel 3D Gaussian (kerbl20233d, ) appearance attributes, such as color and opacity, for novel view rendering, (3) a motion head estimates the time-conditioned movements of Gaussian primitives towards a target timestamp, allowing us to track its temporal evolution.
-Benefiting from the unified architecture, MoVieS can be trained on large-scale datasets featuring both static (zhou2018stereo, ; wang2020tartanair, ; li2023matrixcity, ) and dynamic (mehl2023spring, ; cabon2020virtual, ) scenes, as well as point tracking datasets (zheng2023pointodyssey, ; karaev2023dynamicstereo, ; jin2024stereo4d, ).
-At inference time, MoVieS takes a monocular video, whether depicting a static or dynamic scene, and reconstructs per-pixel 3D Gaussian primitives along with their motion attributes at any target timestamp, enabling novel view synthesis, depth estimation, and 3D point tracking in a single model.
-Extensive experiments on diverse benchmarks (zhou2018stereo, ; yoon2020novel, ; gao2022dynamic, ; koppula2024tapvid, ) show that MoVieS achieves competitive performance across a variety of 4D perception tasks, while being several orders of magnitude faster than the existing state of the art.
+As shown in Figure 1, MoVieS is built upon a large-scale pretrained transformer backbone [47, 68], which encodes each video frame independently and aggregates their information via attentions [66].
+The aggregated features are then processed by specialized prediction heads: (1) a depth head estimates depth for each input frame, (2) a splatter head predicts per-pixel 3D Gaussian [24] appearance attributes, such as color and opacity, for novel view rendering, (3) a motion head estimates the time-conditioned movements of Gaussian primitives towards a target timestamp, allowing us to track its temporal evolution.
+Benefiting from the unified architecture, MoVieS can be trained on large-scale datasets featuring both static [97, 75, 29] and dynamic [44, 3] scenes, as well as point tracking datasets [96, 21, 20].
+During inference, it takes a monocular video, whether depicting a static or dynamic scene, and reconstructs per-pixel 3D Gaussian primitives along with their motion attributes at any target timestamp, enabling view synthesis, depth estimation, and 3D point tracking in a single model.
+Extensive experiments on diverse benchmarks [97, 90, 14, 25] reveal that MoVieS achieves competitive performance across a variety of 4D perception tasks, while being several orders of magnitude faster than the existing state of the art.
 Furthermore, empowered by novel view synthesis as a proxy task, MoVieS enables dense motion understanding from sparse tracking supervision.
-This naturally gives rise to a variety of zero-shot applications, further broadening the potential of our approach.
+This naturally gives rise to various applications, such as scene flow estimation and moving object segmentation, further broadening the potential of our approach.
 In summary, our main contributions include:
-  • We introduce MoVieS, the first feed-forward framework that jointly models appearance, geometry and motion for 4D scene perception from monocular videos.
+  • We introduce MoVieS, a novel feed-forward framework that jointly models appearance, geometry and motion for 4D scene perception from monocular videos.
   • Dynamic splatter pixels are proposed to represent dynamic 3D scenes as renderable deforming 3D particles, bridging novel view synthesis and dynamic geometry reconstruction.
   • MoVieS delivers strong performance and orders of magnitude speedups for 4D reconstruction, and naturally enables a wide range of applications in a zero-shot manner.
-Figure 1: Overview.
-MoVieS consists of a shared image encoder, an attention-based feature backbone (Section 3.2.1), and three heads (Section 3.2.2) that simultaneously predict appearance, geometry and motion.
-Image shortcut for splatter head and time-varying Gaussian attributes are omitted for brevity.
 2 Related Work
+Figure 2: Motion Head.
+Given $M$ query timesteps, the proposed motion head is conditioned via adaptive layer normalization (AdaLN) and predicts 3D displacements for each input pixel.
+After rasterization using the $M$ corresponding query-time cameras, output images in shape $M\times 3\times H\times W$ are rendered for supervision.
+Gaussian attribute deformation $\Delta\mathbf{a}$ is omitted for brevity.
 2.1 Feed-forward 3D Reconstruction
-Most traditional 3D reconstruction methods require dense multi-view supervision and per-scene optimization from scratch at test time (hartley2003multiple, ; davison2007monoslam, ; agarwal2011building, ; mur2015orb, ; schonberger2016structure, ; mildenhall2020nerf, ; kerbl20233d, ).
-Recent learning-based approaches leverage large-scale data priors to enable feed-forward 3D reconstruction that directly regresses geometric structures such as camera poses and depth (wang2024vggsfm, ; wang2025moge, ; ranftl2020towards, ; piccinelli2024unidepth, ; bochkovskii2024depth, ; yang2024depth, ; hu2024metric3d, ).
-With the advent of differentiable rendering techniques like 3D Gaussian Splatting (3DGS) (kerbl20233d, ), several works propose to predict pixel-aligned Gaussian attributes for novel view synthesis (charatan2024pixelsplat, ; chen2024mvsplat, ; zhang2024gs, ; ziwen2024long, ; xu2024depthsplat, ).
-We build on these advances and extend them to dynamic 3DGS representations for novel timestamps.
-DUSt3R (wang2024dust3r, ) pioneered direct regression of pixel-aligned pointmaps in a canonical space from image pairs.
-Subsequent works improved this framework by incorporating feature matching (leroy2024grounding, ), supporting multi-view inputs (wang2025d, ; yang2025fast3r, ; zhang2025flare, ; tang2024mv, ), adapting to streaming video (liu2024slam3r, ; murai2024mast3r, ), or bridging view synthesis with 3DGS (smart2024splatt3r, ; ye2024no, ; zhang2025flare, ; tang2024mv, ).
-VGGT (wang2025vggt, ) further advances this line by using a stronger image encoder (oquab2024dinov2, ) and multiple task-specific heads to predict all key 3D properties from as few as one to hundreds of views.
-We adopt VGGT as the backbone of MoVieS for its strong visual representations.
-However, all these methods are restricted to static scenes and overlook dynamic environments with moving objects—the main focus of our work.
-2.2 Dynamic Geometry Reconstruction
-Compared to static scene reconstruction, which has been studied for decades, dynamic scene reconstruction remains relatively underexplored.
-Some works extend the DUSt3R framework to dynamic settings in a plug-and-play manner (chen2025easi3r, ), or leverage foundation models (zhang2024monst3r, ; lu2024align3r, ; yao2025uni4d, ) by incorporating monocular depth (bochkovskii2024depth, ; piccinelli2025unidepthv2, ), optical flow (teed2020raft, ), or 2D point tracking (karaev2024cotracker, ).
-Dynamic Point Map (sucar2025dynamic, ), a concurrent work, proposes a time-invariant DUSt3R variant by predicting each viewpoint’s pointmap at another timestamp.
-The most related work, DynaDUSt3R (jin2024stereo4d, ), adds pixel-level motion supervision to model dynamic geometry.
-However, all these methods are limited to two-frame inputs and produce only point clouds, without high-quality visual reconstructions.
-To go beyond two views, CUT3R (wang2025continuous, ) uses a recurrent model to incrementally update scene representations from streaming video for static and dynamic scenes.
-Other methods (xu2025geometrycrafter, ; jiang2025geo4d, ) treat dynamic reconstruction as conditional video generation, fine-tuning diffusion models to estimate pointmaps probabilistically.
-However, these require multiple network passes per video.
-2.3 Dynamic Novel View Synthesis
-Instead of geometric structure, dynamic novel view synthesis emphasizes visual results and view-dependent effects.
-Early attempts utilize neural radiance field (mildenhall2020nerf, ; fridovich2023k, ; cao2023hexplane, ) to represent dynamic scenes, either from multi-view videos (bansal20204d, ; wang2022fourier, ; li2022neural, ), or monocular ones (pumarola2021d, ; park2021nerfies, ; li2023dynibar, ; zhao2024pgdvs, ).
-3DGS (kerbl20233d, ), as a more efficient renderable representation, has been applied for dynamic view synthesis using 4D primitives with a time dimension (yang2024real, ; li2024spacetime, ) or an additional deformable field (wu20244d, ; luiten2024dynamic, ; yang2024deformable, ; lin2024gaussian, ; liu2025modgs, ).
-Leveraging the low-dimensional nature of scene motion, several works (wang2024shape, ; sun2024splatter, ; stearns2024dynamic, ; lei2025mosca, ; lin2025omniphysgs, ) express motion explicitly and learn Gaussian trajectories across monocular frames.
-However, these methods are trained from scratch and require iterative optimization for scene adaptation.
-Off-the-shelf point tracking (doersch2023tapir, ; karaev2024cotracker, ) or optical flow (teed2020raft, ) models are also essential to provide supervision signals.
-The most related works to ours are BTimer (liang2024feed, ) and NutWorld (shen2025seeing, ), both of which utilize a feed-forward approach to estimate 3DGS attributes from monocular videos.
-However, BTimer predicts independent Gaussian chunks per timestamp without modeling frame relations and needs an enhancer module for smooth intermediate frames.
-NutWorld models Gaussian motion but lacks explicit supervision, relying heavily on pretrained depth (chen2025video, ) and flow (xu2023unifying, ), and uses an orthographic camera, which may further lead to projection distortions.
+Traditional 3D reconstruction relies on dense multi-view supervision and per-scene optimization [16, 46, 56, 45].
+Recent learning-based methods leverage large-scale priors to enable feed-forward prediction of depth, pose, and other 3D properties [69, 54, 51, 17].
+With differentiable rendering techniques such as 3D Gaussian Splatting (3DGS) [24], some works further regress pixel-wise Gaussian attributes for novel view synthesis [5, 9, 92].
+DUSt3R [74] pioneered the direct regression of pixel-aligned pointmaps in a canonical space from image pairs.
+Its extensions [27, 67, 39, 59, 94] generalize this framework with multi-view, streaming, and 3DGS integration.
+VGGT [68] unifies these advances with a strong image encoder [47] and task-specific heads.
+However, all of these works are limited to static scenes.
+In contrast, in this work, we extend feed-forward 3DGS reconstruction to dynamic scenarios with moving objects.
+2.2 Dynamic Reconstruction and View Synthesis
+Compared to static reconstruction, dynamic reconstruction remains underexplored.
+Extensions of DUSt3R [8, 61, 20] handle dynamic reconstruction in a plug-and-play manner [8] or leveraging foundation models [91, 42, 87] with monocular depth [2, 50], optical flow [64], and point tracking [22], but remain limited to two-frame inputs and output only sparse point clouds.
+CUT3R [73] extends to video inputs via recurrent updates, while diffusion-based methods [82, 19] treat reconstruction as conditional generation, though at the cost of multiple passes per sequence.
+Instead of geometric structure, dynamic novel view synthesis focuses on rendering quality and view-dependent effects.
+NeRF-based methods [45, 12, 4] reconstruct dynamic scenes from multi-view [1, 70, 28] or monocular videos [52, 48, 32, 95], while 3DGS [24] extends this with 4D primitives [85, 30] or deformable fields [76, 43, 86, 35, 37].
+Recent works further express Gaussian motion explicitly [72, 62, 60, 26, 36], but require scratch training, iterative optimization, and external supervision from point tracking or optical flow estimation [11, 64].
+For feed-forward 4D view synthesis, STORM [83] is designed for outdoor driving scenes from multi-view videos.
+BTimer [33] and NutWorld [57] estimate 3DGS attributes from monocular inputs.
+However, BTimer predicts independent pixel-wise 3D Gaussians per timestamp without modeling their frame-wise relations and needs an extra enhancer module for smooth intermediate frames.
+NutWorld models Gaussian motion but lacks explicit supervision, relying heavily on pretrained depth [6] and flow [80] estimation models, and uses an orthographic camera, which further leads to projection distortions.
 3 Method
-Following previous studies on monocular 4D reconstruction (zhao2024pgdvs, ; wang2024shape, ; liu2025modgs, ; lei2025mosca, ; liang2024feed, ), given a posed video $\mathcal{V}\coloneqq\{\mathbf{I}_{i},\mathbf{P}_{i},\mathbf{K}_{i}\}_{i=1}^{N}$ with frames $\mathbf{I}_{i}\in\mathbb{R}^{3\times H\times W}$, camera poses $\mathbf{P}_{i}\in\mathbb{SE}(3)$, intrinsics $\mathbf{K}_{i}\in\mathbb{R}^{3\times 3}$ and timestamps $t_{i}\in[0,1]$, we aim to train a generic model that jointly reconstruct the underlying appearance, geometry and motion of dynamic scenes.
+Following previous studies on monocular 4D reconstruction [95, 72, 37, 26, 33, 83], given a posed video $\mathcal{V}\coloneqq\{\mathbf{I}_{i},\mathbf{P}_{i},\mathbf{K}_{i},t_{i}\}_{i=1}^{N}$ with frames $\mathbf{I}_{i}\in\mathbb{R}^{3\times H\times W}$, camera poses $\mathbf{P}_{i}\in\mathbb{SE}(3)$, intrinsics $\mathbf{K}_{i}\in\mathbb{R}^{3\times 3}$ and timestamps $t_{i}\in[0,1]$, we aim to train a generic model that jointly reconstructs its appearance, geometry, and motion.
+We leave the handling of unposed videos to future work.
 3.1 Dynamic Splatter Pixel
 To model 3D scenes with moving contents, we propose a novel representation, namely dynamic splatter pixel, which decomposes dynamic scenes into a set of static Gaussian primitives and their corresponding deformation fields.
-Given an input video $\mathcal{V}$, each pixel in the $i$-th frame $\mathbf{I}_{i}$ is associated with a splatter pixel $\mathbf{g}_{i}$ (charatan2024pixelsplat, ; szymanowicz2024splatter, ; zhang2024gs, ) in a shared canonical space of the first frame camera’s coordinate system.
-Each $\mathbf{g}\coloneqq\{\mathbf{x},\mathbf{a}\}$ is parameterized by its position $\mathbf{x}\in\mathbb{R}^{3}$ in the canonical space and other rendering attributes $\mathbf{a}\in\mathbb{R}^{11}$, including rotation quaternion $\mathbf{q}\in\mathbb{R}^{4}$, scale $\mathbf{s}\in\mathbb{R}^{3}$, opacity $\alpha\in\mathbb{R}$, and color $\mathbf{c}_{\text{rgb}}\in\mathbb{R}^{3}$ (kerbl20233d, ).
+Given an input video $\mathcal{V}$, each pixel in the $i$-th frame $\mathbf{I}_{i}$ is associated with a splatter pixel $\mathbf{g}_{i}$ [5, 63, 92] in a shared canonical space of the first frame camera’s coordinate system.
+Each $\mathbf{g}\coloneqq\{\mathbf{x},\mathbf{a}\}$ is parameterized by its position $\mathbf{x}\in\mathbb{R}^{3}$ in the canonical space and other rendering attributes $\mathbf{a}\in\mathbb{R}^{11}$, including rotation quaternion $\mathbf{q}\in\mathbb{R}^{4}$, scale $\mathbf{s}\in\mathbb{R}^{3}$, opacity $\alpha\in\mathbb{R}$, and color $\mathbf{c}_{\text{rgb}}\in\mathbb{R}^{3}$ [24].
 Considering splatter pixels were originally designed for static scenes, we decouple motion from geometric structure to adapt them for dynamic scenes.
@@
 By combining static splatter pixels and their deformation fields, we establish the correspondence between each Gaussian primitive and its temporal dynamics, thereby enabling dynamic scene modeling and dense motion estimation.
-In practice, we find that making only the scale and rotation attributes time-dependent, i.e., $\Delta\mathbf{a}(t)\in\mathbb{R}^{7}$, is sufficient for representing dynamic splatter pixels.
-Figure 2: Motion Head.
-Given $t_{q}$ target timesteps, the proposed motion head is conditioned via adaptive layer normalization (AdaLN) and predicts 3D displacements for each input pixel.
-After rasterization using the $M$ corresponding query-time cameras, output images in shape $M\times 3\times H\times W$ are rendered for supervision.
-Gaussian attribute deformation $\Delta\mathbf{a}$ is omitted for brevity.
-3.2 MoVieS: Unify Appearance, Geometry and Motion
-As illustrated in Figure 1, the framework of the proposed MoVieS provides a unified approach to simultaneously model appearance, geometry and motion of dynamic scenes.
-It consists of a feature backbone with camera and time information for extracting features from input video frames, followed by dedicated heads for depth, splatter and motion estimation.
+3.2 Unify Appearance, Geometry and Motion
 3.2.1 Feature Backbone
-Given an input posed video $\mathcal{V}=\{\mathbf{I}_{i},\mathbf{P}_{i},\mathbf{K}_{i},t_{i}\}_{i=1}^{N}$, we first patchify the input images $\mathbf{I}_{i}$ and use a pretrained image encoder (oquab2024dinov2, ) to extract their features.
-To effectively incorporate camera information, we adopt two complementary strategies to embed the camera parameters into image features: (1) Plücker embedding: camera pose $\mathbf{P}_{i}$ and intrinsics $\mathbf{K}_{i}$ are encoded into pixel-aligned Plücker embeddings (sitzmann2021light, ), which are then downsampled and fused with image features via spatial-wise addition; (2) Camera token: both $\mathbf{P}_{i}$ and $\mathbf{K}_{i}$ are passed through a linear layer to generate a camera token, which is appended to the sequence of image tokens.
-Plücker embedding provides a dense and spatially aligned encoding of camera geometry, while the camera token interacts globally with image tokens through attention, enabling more holistic feature reasoning.
-Ablation study on these two camera injection manners is validated in Section 4.4.
-To inform the model that input images originate from a temporally ordered video, we additionally encode each timestamp $t_{i}\in[0,1]$ using sinusoidal positional encoding (mildenhall2020nerf, ) to produce a time token, which is then concatenated with the aforementioned image and camera tokens.
-After tokenizing input images, camera parameters, and timestamps, we apply geometrically pretrained attention blocks (vaswani2017attention, ) from VGGT (wang2025vggt, ) to enable interactions among image tokens across video frames.
+Given a posed video $\mathcal{V}=\{\mathbf{I}_{i},\mathbf{P}_{i},\mathbf{K}_{i},t_{i}\}_{i=1}^{N}$, we first patchify each input image $\mathbf{I}_{i}$ and use a pretrained image encoder [47] to extract their features as shown in Figure 1.
+To effectively incorporate camera information, we adopt two complementary strategies to embed the camera parameters into image features: (1) Plücker embedding: camera pose $\mathbf{P}_{i}$ and intrinsics $\mathbf{K}_{i}$ are transformed into pixel-aligned Plücker embeddings [58], which are then downsampled and fused with image features via spatial-wise addition; (2) Camera token: $\mathbf{P}_{i}$ and $\mathbf{K}_{i}$ are passed through a linear layer and encoded to a camera token, which is appended to the sequence of image tokens.
+Ablation study on these two camera injection manners is validated in Sec.
+4.4.
+To inform the model that input images originate from a temporally ordered video, we additionally encode the timestamp $t_{i}\in[0,1]$ of input frames by sinusoidal positional encoding [45] to produce a timestamp token, which is then concatenated with the aforementioned image and camera tokens as shown in Figure 1.
+After tokenizing input images, camera parameters, and timestamps, we apply the geometrically pretrained attention blocks [66] from VGGT [68] to enable interactions among image tokens across video frames.
 This produces a set of shared feature tokens enriched with inter-frame context as well as camera and temporal information, which are then used for predicting various properties of dynamic scenes.
+The effect of VGGT initialization is discussed in Sec.
+4.4.
 3.2.2 Prediction Heads
-Shared aggregated video tokens from the feature backbone are fed into three parallel prediction heads that estimate the appearance, geometry, and motion of the dynamic scene, respectively.
-Each head adopts a DPT-style architecture (ranftl2021vision, ) to convert image tokens into dense predictions matching the input resolution, thereby producing dynamic splatter pixels.
+Tokens from the feature backbone are fed into three parallel prediction heads that estimate the appearance, geometry and motion of a dynamic scene respectively, as shown in Figure 1.
+Each head adopts a DPT-style architecture [53] to convert aggregated tokens into dense predictions matching the input resolution and product dynamic splatter pixels.
 Depth and Splatter Head
-Different from previous feed-forward 3DGS reconstruction methods (charatan2024pixelsplat, ; chen2024mvsplat, ; zhang2024gs, ) that use a single head to predict all splatter pixel attributes, we adopt a decoupled design to better leverage geometric priors from the pretrained VGGT (wang2025vggt, ).
-Specifically, a dedicated depth head, initialized from VGGT, is used for geometry prediction to provide spatial grounding for splatter pixel construction, while a separate DPT as splatter head is trained from scratch for appearance rendering.
-We further incorporate a direct RGB shortcut (ye2024no, ) from the input image to the final convolution layer of the splatter head to preserve high-frequency details and enhance color fidelity.
+Different from previous feed-forward 3DGS reconstruction methods [5, 9, 92] that use a single head to predict all splatter pixel attributes, we adopt a decoupled design to better leverage geometric priors from the pretrained VGGT [68].
+Specifically, a dedicated depth head, initialized from VGGT, is used for geometry prediction to provide spatial grounding for splatter pixel construction, while another DPT as splatter head is trained from scratch for appearance rendering.
+We further incorporate a direct RGB shortcut [88] from the input image to the final convolution layer of the splatter head to preserve high-frequency details and enhance color fidelity.
 Motion Head
 To capture scene dynamics, a novel motion head as shown in Figure 2 is introduced to predict dense deformation $\mathbf{m}(t)$ for each dynamic splatter pixel at any target moment.
-Temporal variation is enabled by injecting the sinusoidally encoded query time $t_{q}$ into the image tokens via adaptive layer normalization (xu2019understanding, ) before applying DPT convolutions.
-For each input frame $\mathbf{I}_{i}$ at time $t_{i}$, the motion head predicts its 3D movements toward $t_{q}$ in a shared world coordinate system.
-Thus, for $N$ input frames and $M$ query timestamps, this yields a motion map of shape $M\times N\times 3\times H\times W$.
+Temporal variation is enabled by injecting the sinusoidally encoded query time $t_{q}$ into the aggregated tokens via adaptive layer normalization [81] before applying DPT convolutions.
+For each input frame $\mathbf{I}_{i}$ at time $t_{i}$, the motion head predicts its 3D movements $\Delta\mathbf{x}$ and Gaussian attribute deformation $\Delta\mathbf{a}$ toward $t_{q}$ in a shared world coordinate system.
+To visualize motion maps, the XYZ coordinates in 3D space are jointly normalized by their shared minimum and maximum values to the range $[0,1]$, and then mapped to RGB channels for visualization.
+Table 1: Training Datasets. Eight datasets from diverse sources are utilized to train MoVieS at scale.
+Dataset Dynamic? Depth? Tracking? Real? #Scenes
+RealEstate10K [97] 70K
+TartanAir [75] 0.4K
+MatrixCity [29] 4.5K
+PointOdyssey [96] 0.1K
+DynamicReplica [21] 0.5K
+Spring [44] 0.03K
+VKITTI2 [3] 0.1K
+Stereo4D [20] 98K
+Figure 3: Novel View Synthesis for Dynamic Scenes.
+Given a monocular video, we compare synthesized novel views of different methods.
+Invisible regions are rendered as black or white, depending on the implementation.
+More results are in the supplementary material.
 3.3 Training
-Table 1: Training Datasets. Nine datasets from diverse sources are utilized to train MoVieS at scale. “#Repeat” denotes dataset duplication count during integration to balance their contributions.
-Dataset Dynamic? Depth? Tracking? Real? #Scenes #Frames #Repeat
-RealEstate10K (zhou2018stereo, ) 70K 6.36M 1$\times$
-TartanAir (wang2020tartanair, ) 0.4K 0.49M 100$\times$
-MatrixCity (li2023matrixcity, ) 4.5K 0.31M 10$\times$
-PointOdyssey (zheng2023pointodyssey, ) 0.1K 0.18M 1000$\times$
-DynamicReplica karaev2023dynamicstereo 0.5K 0.26M 100$\times$
-Spring (mehl2023spring, ) 0.03K 0.003M 2000$\times$
-VKITTI2 (cabon2020virtual, ) 0.1K 0.03M 500$\times$
-Stereo4D (jin2024stereo4d, ) 98K 19.6M 1$\times$
 3.3.1 Dataset Curation
-An ideal dataset for dynamic scene reconstruction would include synchronized multi-view videos with dense depth and point tracking annotations.
+An ideal dataset for dynamic reconstruction would include synchronized multi-view videos with dense depth and point tracking annotations.
 However, such data is infeasible to capture and annotate at scale in practice.
-Instead, we leverage a diverse set of open-source datasets (zhou2018stereo, ; wang2020tartanair, ; li2023matrixcity, ; mehl2023spring, ; cabon2020virtual, ; zheng2023pointodyssey, ; karaev2023dynamicstereo, ; jin2024stereo4d, ), each providing complementary supervision, as shown in Table 1.
-With the flexible model design, MoVieS can be jointly trained on these heterogeneous sources by aligning objectives to their respective annotations.
-More details about data curation are provided in Appendix A.
+Instead, we leverage a diverse set of open-source datasets [97, 75, 29, 44, 3, 96, 21, 20], each providing complementary supervision, as shown in Table 1.
+With the flexible model design, MoVieS can be trained on these heterogeneous sources by aligning objectives to their respective annotations.
+For datasets without an official test split, we randomly sample 10% of the data as the test set for generalization evaluation.
+Details about data curation are in the supplementary material.
 3.3.2 Objectives
@@
 Depth and Rendering Losses
-Depth loss is computed as the mean squared error (MSE) between the predicted and ground truth depth maps, along with their spatial gradients, after filtering out invalid values.
-Rendering loss combines pixel-wise MSE and perceptual loss (zhang2018perceptual, ) between 3DGS-rendered images under corresponding camera views and the video frames at target timestamps.
+Depth loss is computed as the mean squared error (MSE) between the predicted depth maps $\hat{D}_{i}$ and ground truth depth maps $D_{i}$, along with their spatial gradients, after filtering out invalid values.
+$\mathcal{L}_{\text{depth}}\coloneq\sum_{i=1}^{N}\|D_{i}-\hat{D}_{i}\|_{2}+\|\nabla D_{i}-\nabla\hat{D}_{i}\|_{1}.$ (3)
+Rendering loss combines pixel-wise MSE and perceptual loss [93] between 3DGS-rendered images $\hat{I}$ under $M$ camera views and the video frames at corresponding target timestamps, which are randomly sampled during training.
+$\mathcal{L}_{\text{rendering}}\coloneq\sum_{v=1}^{M}\|I_{v}-\hat{I}_{v}\|_{2}+\lambda_{\text{LPIPS}}\cdot\text{LPIPS}(I_{v},\hat{I}_{v}).$ (4)
+Weighting terms $\lambda_{\text{d}}$, $\lambda_{\text{r}}$ and $\lambda_{\text{LPIPS}}$ are set to $1$, 11 and $0.5$ by default respectively.
 Motion Loss
@@
 $\mathcal{L}_{\text{motion}}$ $\coloneqq\lambda_{\text{pt}}\mathcal{L}_{\text{pt}}+\lambda_{\text{dist}}\mathcal{L}_{\text{dist}}$ (5)
-$\ =\frac{1}{P}\sum_{i\in\Omega}\lambda_{\text{pt}}\|\Delta\hat{\mathbf{x}}_{i}-\Delta\mathbf{x}_{i}\|_{1}+\frac{1}{P^{2}}\sum_{(i,j)\in\Omega\times\Omega}\lambda_{\text{dist}}\|\Delta\hat{\mathbf{x}}_{i}\cdot\Delta\hat{\mathbf{x}}_{j}^{\top}-\Delta\mathbf{x}_{i}\cdot\Delta\mathbf{x}_{j}^{\top}\|_{1},$ (4)
-where $\Omega$ denotes the set of all valid $P$ tracked points, and $\Omega\times\Omega$ is its Cartesian product.
-Ablation study on the effectiveness of these two types of motion supervision is presented in Section 4.4.
+$\ =\frac{1}{P}\sum_{i\in\Omega}\lambda_{\text{pt}}\|\Delta\hat{\mathbf{x}}_{i}-\Delta\mathbf{x}_{i}\|_{1}$ (6)
+$\ +\frac{1}{P^{2}}\sum_{(i,j)\in\Omega\times\Omega}\lambda_{\text{dist}}\|\Delta\hat{\mathbf{x}}_{i}\cdot\Delta\hat{\mathbf{x}}_{j}^{\top}-\Delta\mathbf{x}_{i}\cdot\Delta\mathbf{x}_{j}^{\top}\|_{1},$ (7)
+where $\Omega$ denotes the set of valid $P$ tracked points, and $\Omega\times\Omega$ is its Cartesian product.
+Weighting term $\lambda_{\text{m}}$, $\lambda_{\text{pt}}$, and $\lambda_{\text{dist}}$ are set to $10$, 11, and $10$ respectively.
+Ablation study on the effectiveness of these two types of motion supervision is presented in Sec.
+4.4.
 Normalization
-Similar to VGGT (wang2025vggt, ), we normalize the 3D scene scale by the average Euclidean distance from each 3D point to the origin of the canonical world coordinate system.
-As a result, unlike some other reconstruction methods (wang2024dust3r, ; leroy2024grounding, ; wang2025continuous, ), we do not apply additional normalization in the depth or motion loss.
-We also omit confidence-aware weighting (wang2024dust3r, ; wang2025vggt, ) for simplicity.
+Similar to VGGT [68], we normalize the 3D scene scale by the average Euclidean distance from each 3D point to the origin of the canonical world coordinate system.
+As a result, unlike some other reconstruction methods [74, 27, 73], we do not apply additional normalization in the depth or motion loss.
+We also omit confidence-aware weighting [74, 68] for simplicity and more stable training.
 4 Experiments
-4.1 Experimental Settings
-Implementation
-MoVieS is built on a geometrically pretrained transformer, VGGT (wang2025vggt, ), with splatter head and camera/time embeddings trained from scratch.
-AdamW (loshchilov2017decoupled, ) with cosine learning rate scheduling and linear warm-up is used for optimization.
-We observed that training MoVieS is particularly unstable, likely arising from sparse annotations and the heterogeneous nature of the training data.
-A curriculum strategy is employed that gradually increases the training complexity, involving (1) pretraining on static scenes, (2) dynamic scenes with varying views, and (3) fine-tuning on high resolution.
-Several techniques, such as gsplat (ye2025gsplat, ) rendering backend, DeepSpeed (rasley2020deepspeed, ), gradient checkpointing (chen2016training, ), gradient accumulation, and bf16 mixed precision, are adopted to improve memory and computation efficiency.
-Training completes in approximately 5 days using 32 H20 GPUs.
-Please refer to Appendix A for more details about implementations.
-Evaluation
-We evaluate the capabilities of MoVieS on two primary tasks: novel view synthesis (Section 4.2) and 3D point tracking (Section 4.3).
-Following prior works (charatan2024pixelsplat, ; chen2024mvsplat, ; xu2024depthsplat, ; wang2024shape, ; sun2024splatter, ; lei2025mosca, ; liang2024feed, ), RealEstate10K (zhou2018stereo, ) is used to evaluate novel view synthesis performance on static scenes.
-DyCheck (gao2022dynamic, ) and NVIDIA dynamic scene dataset (yoon2020novel, ) are adopted for dynamic scenes, in terms of PSNR, SSIM and LPIPS (zhang2018perceptual, ).
-Since the DyCheck dataset contains invisible regions in the target novel views, we compute reconstruction metrics (denoted by a prefix “m”) using the provided covisibility masks, ensuring a fair comparison across methods.
-For 3D point tracking, the TAPVid-3D (koppula2024tapvid, ) benchmark serves as the evaluation protocol, covering both indoor and outdoor scenes across three datasets.
-End-point Euclidean error in the 3D space (EPE3D) and the percentage of 3D points within $0.05$ and $0.10$ units of the ground-truth locations ($\delta^{0.05}_{3D}$ and $\delta^{0.10}_{3D}$) are utilized as metrics.
-4.2 Novel View Synthesis
-4.2.1 Static Scene View Synthesis
-As a special case of dynamic scenes, we first evaluate MoVieS on a static dataset RealEstate10K (zhou2018stereo, ), comparing against several state-of-the-art feed-forward static scene reconstruction methods, including DepthSplat (xu2024depthsplat, ), our reimplementation of GS-LRM (zhang2024gs, ), and our method pretrained solely on the first-stage static reconstruction.
-As shown in Table 2, although MoVieS is primarily designed for dynamic scenes, it maintains competitive performance on static scenes.
-Notably, when processing static inputs, our predicted motion naturally converges to zero, demonstrating the ability of MoVieS to implicitly differentiate between static and dynamic regions without explicit supervision.
-4.2.2 Dynamic Scene View Synthesis
-Dynamic scene view synthesis is evaluated on two benchmarks: (1) DyCheck (gao2022dynamic, ), with 7 daily videos captured by 3 synchronized cameras, and (2) NVIDIA (yoon2020novel, ), containing 9 dynamic scenes recorded by a 12-camera static rig.
-For DyCheck, we sample 13 frames of size $518\times 518$ from a random clip from iPhone-record videos, and include the other two cameras for evaluation.
-For NVIDIA, 12 input views at $379\times 672$ resolution are sampled in a round-robin style (gao2021dynamic, ; liu2023robust, ; liang2024feed, ), where the i-th frame comes from the i-th camera at timestep i.
-We compare against three state-of-the-art optimization-based methods (wang2024shape, ; sun2024splatter, ; lei2025mosca, ) for dynamic 3D Gaussian reconstruction, and also report static feed-forward baselines for reference.
-As shown in Table 2, MoVieS achieves competitive or superior performance compared to baselines, while requiring only 0.93s per scene, which is orders of magnitude faster than prior approaches that rely on heavy pretrained models and complex multi-stage pipelines.
-Qualitative visualizations in Figure 3 further highlight the strength of our approach.
-While MoSca (lei2025mosca, ) demonstrates impressive performance, it struggles with sparse inputs and often overfits to seen poses, producing spiky and over-saturated artifacts under novel views and timesteps.
-In contrast, MoVieS leverages large-scale learned priors to generalize more effectively, yielding smoother and more realistic results.
-To ensure fair comparison and better reflect real-world scenarios, no video masks for dynamic objects are used in our experiments.
-It poses a major challenge for optimization-based methods like Shape-of-Motion (wang2024shape, ) and Splatter-a-Video (sun2024splatter, ), which rely heavily on explicit motion segmentation.
-The challenge is particularly evident on the NVIDIA dataset, where significant camera shake hinders disentangling scene dynamics, leading to notably degraded performance and even worse than static baselines.
-In contrast, our method exhibits strong robustness by directly learning to model motion.
-Figure 3: Novel View Synthesis for Dynamic Scenes.
-Given a monocular video, we compare synthesized views from a novel view across different methods.
-Regions invisible in the input are rendered as black or white, depending on the rendering implementation.
-More results in Figure 6.
-Table 2: Evaluation on Novel View Synthesis. The best, second best and third best results are highlighted for clarity. $\dagger$ indicates our reimplemented version of GS-LRM (zhang2024gs, ).
+Table 2: Evaluation on Novel View Synthesis.
+The best, second best and third best results are highlighted for clarity.
+$\dagger$ indicates our reimplemented version of GS-LRM [92].
+“Ours (static)” refers to our method pretrained solely on static datasets without the motion head.
+The same camera parameters are provided for all methods for fair comparison.
 Novel View Synthesis Time Per Scene RealEstate10K DyCheck NVIDIA
@@
 Static Feed-forward
-DepthSplat (xu2024depthsplat, ) 0.60s 26.57 80.06 0.124 13.83 0.4364 0.3850 17.16 0.5002 0.3023
-GS-LRM (zhang2024gs, ) 0.57s 26.94 79.13 0.139 14.60 0.4535 0.3775 17.83 0.4988 0.3265
-Ours (static) 0.84s 27.60 81.25 0.113 15.24 0.4784 0.3783 18.73 0.5042 0.2959
+DepthSplat [79] 0.60s 26.57 80.06 0.124 13.83 43.64 0.3850 17.16 50.02 0.3023
+GS-LRM [92] 0.57s 26.94 79.13 0.139 14.60 45.35 0.3775 17.83 49.88 0.3265
+Ours (static) 0.84s 27.60 81.25 0.113 15.24 47.84 0.3783 18.73 50.42 0.2959
 Optimization-based
-Splatter-a-Video (sun2024splatter, ) 37min - - - 13.61 0.3131 0.5706 14.39 0.2538 0.5983
-Shape-of-Motion (wang2024shape, ) 10min - - - 17.96 0.5662 0.3463 15.30 0.3169 0.5087
-MoSca (lei2025mosca, ) 45min - - - 18.24 0.5514 0.3698 21.45 0.7123 0.2653
-Ours 0.93s 26.98 81.75 0.111 18.46 0.5887 0.3094 19.16 0.5141 0.3152
-4.3 3D Point Tracking
-Trained on large-scale point tracking datasets (zheng2023pointodyssey, ; karaev2023dynamicstereo, ; jin2024stereo4d, ), the proposed method can also densely track any 3D point corresponding to a pixel across video frames (see Figure 7 in the appendix).
-We compare MoVieS against three strong baselines, including two state-of-the-art 2D point tracking methods, BootsTAP (doersch2024bootstap, ) and CoTracker3 (karaev2024cotracker, ), as well as a native 3D point tracking approach, SpatialTracker (xiao2024spatialtracker, ).
-For 2D trackers, a recent video depth estimation model (chen2025video, ) and ground-truth camera intrinsics are used to unproject tracked points into 3D space.
-To account for scale differences across methods, we normalize all predicted 3D points by their median norm before evaluation.
-Quantitative results are reported in Table 3.
-While 3D-based SpatialTracker generally outperforms 2D-based approaches, all of them rely heavily on pretrained monocular depth estimators for geometry reasoning, introducing significant noise and inconsistency in the 3D space.
-In contrast, MoVieS directly estimates 3D point positions in a shared world coordinate, enabling more accurate and robust 3D tracking, and achieves consistently superior or competitive performance across all datasets.
+Splatter-a-Video [62] 37min - - - 13.61 31.31 0.5706 14.39 25.38 0.5983
+Shape-of-Motion [72] 10min - - - 17.96 56.62 0.3463 15.30 31.69 0.5087
+MoSca [26] 45min - - - 18.24 55.14 0.3698 21.45 71.23 0.2653
+Ours 0.93s 26.98 81.75 0.111 18.46 58.87 0.3094 19.16 51.41 0.3152
 Table 3: Evaluation on 3D Point Tracking.
 The best, second best and third best results are highlighted for clarity.
-$\dagger$ denotes 3D points obtained by combining a depth estimation model (chen2025video, ).
+$\dagger$ denotes combining a depth estimation model [6].
+All methods are given the same camera information for unprojecting 3D tracked points.
 3D Point Tracking Aria Digital Twin DriveTrack Panoptic Studio
-$\downarrow$ EPE3D $\uparrow$ $\delta^{0.05}_{3D}$uparrow $\delta^{0.10}_{3D}$downarrow EPE3D $\uparrow$ $\delta^{0.05}_{3D}$uparrow $\delta^{0.10}_{3D}$downarrow EPE3D $\uparrow$ $\delta^{0.05}_{3D}$uparrow $\delta^{0.10}_{3D}$
-BootsTAPIR (doersch2024bootstap, ) 0.5539 17.73% 32.97% 0.0617 55.82% 75.66% 0.0650 69.28% 87.95%
-CoTracker3 (karaev2024cotracker, ) 0.5614 19.88% 35.82% 0.0637 55.30% 77.55% 0.0617 69.27% 88.04%
-SpatialTracker (xiao2024spatialtracker, ) 0.5413 18.08% 38.23% 0.0648 56.58% 80.67% 0.0519 72.91% 89.86%
+$\downarrow$ EPE3D $\uparrow$ $\delta^{0.05}_{3D}$ $\uparrow$ $\delta^{0.10}_{3D}$ $\downarrow$ EPE3D $\uparrow$ $\delta^{0.05}_{3D}$ $\uparrow$ $\delta^{0.10}_{3D}$ $\downarrow$ EPE3D $\uparrow$ $\delta^{0.05}_{3D}$ $\uparrow$ $\delta^{0.10}_{3D}$
+BootsTAPIR [10] 0.5539 17.73% 32.97% 0.0617 55.82% 75.66% 0.0650 69.28% 87.95%
+CoTracker3 [22] 0.5614 19.88% 35.82% 0.0637 55.30% 77.55% 0.0617 69.27% 88.04%
+SpatialTracker [77] 0.5413 18.08% 38.23% 0.0648 56.58% 80.67% 0.0519 72.91% 89.86%
 Ours 0.2153 52.05% 71.63% 0.0472 60.63% 79.87% 0.0352 87.88% 94.61%
-4.4 Ablation and Analysis
 Figure 4: Motion Visualization for Ablation Studies.
@@
 + Camera token 26.81 78.65 0.121
-Ours (static) 27.60 81.25 0.113
+Ours 27.60 81.25 0.113
 Table 5: Ablation study on motion supervision.
 Motion Supervision Aria Digital Twin
-$\downarrow$ EPE3D $\uparrow$ $\delta^{0.05}_{3D}$uparrow $\delta^{0.10}_{3D}$
+$\downarrow$ EPE3D $\uparrow$ $\delta^{0.05}_{3D}$ $\uparrow$ $\delta^{0.10}_{3D}$
 N/A 0.7938 19.58% 32.86%
@@
 Motion & View Synthesis DyCheck NVIDIA Aria Digital Twin
-$\uparrow$ mPSNR $\uparrow$ mSSIM $\downarrow$ mLPIPS $\uparrow$ PSNR $\uparrow$ SSIM $\downarrow$ LPIPS $\downarrow$ EPE3D $\uparrow$ $\delta^{0.05}_{3D}$uparrow $\delta^{0.10}_{3D}$
-NVS w/o motion 15.82 0.4796 0.3741 18.38 0.4856 0.3010 0.7938 19.58% 32.86%
-Motion w/o NVS 16.26 0.4556 0.3461 18.98 0.4939 0.3207 0.3801 24.72% 42.92%
-Ours 18.46 0.5887 0.3094 19.16 0.5041 0.3152 0.2153 52.05% 71.63%
+$\uparrow$ mPSNR $\uparrow$ mSSIM $\downarrow$ mLPIPS $\uparrow$ PSNR $\uparrow$ SSIM $\downarrow$ LPIPS $\downarrow$ EPE3D $\uparrow$ $\delta^{0.05}_{3D}$ $\uparrow$ $\delta^{0.10}_{3D}$
+NVS w/o motion 15.82 47.96 0.3741 18.38 48.56 0.3010 0.7938 19.58% 32.86%
+Motion w/o NVS 16.26 45.56 0.3461 18.98 49.39 0.3207 0.3801 24.72% 42.92%
+Ours 18.46 58.87 0.3094 19.16 50.41 0.3152 0.2153 52.05% 71.63%
+Figure 5: Zero-shot Applications.
+Predicted motion maps from our model can be directly applied to downstream tasks, such as (a) scene flow estimation and (b) moving object segmentation, in a zero-shot manner, without any task-specific fine-tuning or supervision.
+4.1 Experimental Settings
+Implementation
+MoVieS is built on a geometrically pretrained transformer, VGGT [68], via full fine-tuning.
+The splatter head and camera/time embeddings are trained from scratch.
+These new embeddings are zero-initialized to provide a stable initialization without interfering with the pretrained model.
+AdamW [41] with cosine learning rate scheduling and linear warm-up is used for optimization.
+We observed that training MoVieS is particularly unstable, likely due to sparse annotations and the heterogeneous nature of training datasets.
+A curriculum strategy is used that gradually increases the training complexity, involving (1) pretraining on static scenes, (2) dynamic scenes with varying views, and (3) fine-tuning in high resolution.
+Several techniques, such as gsplat rendering backend [89], DeepSpeed [55], gradient checkpointing [7], gradient accumulation, and bf16 mixed precision, are adopted to improve memory and computation efficiency.
+Training completes in approximately 5 days using 32 H20 GPUs.
+More details are provided in the supplementary material.
+Evaluation
+We evaluate MoVieS on two primary tasks: novel view synthesis (Sec.
+4.2) and 3D point tracking (Sec.
+4.3).
+Considering pose-aware scenarios, we provide the same camera parameters for all methods to ensure a fair comparison.
+Following prior works [5, 9, 79, 72, 62, 26, 33], RealEstate10K [97] is used to evaluate novel view synthesis performance on static scenes.
+DyCheck [14] and NVIDIA dynamic scene dataset [90] are adopted for dynamic scenes, in terms of PSNR, SSIM and LPIPS [93].
+Since the DyCheck dataset contains invisible regions in the target novel views, we compute reconstruction metrics using the provided covisibility masks (denoted by a prefix “m”), ensuring a fair comparison across methods.
+For 3D point tracking, the TAPVid-3D [25] benchmark serves as the evaluation protocol, covering both indoor and outdoor scenes across three datasets.
+End-point Euclidean error in the 3D space (EPE3D) and the percentage of 3D points within $0.05$ and $0.10$ units of the ground-truth locations ($\delta^{0.05}_{3D}$ and $\delta^{0.10}_{3D}$) are utilized as metrics.
+A large pretrained video depth estimation model, Video Depth Anything [6], and the same camera information are combined to unproject the predicted 2D tracking points of baselines into the 3D space for fair comparison.
+4.2 Novel View Synthesis
+4.2.1 Static Scene View Synthesis
+As a special case of dynamic scenes, we first evaluate MoVieS on a static dataset, RealEstate10K [97], comparing against several state-of-the-art feed-forward static scene reconstruction methods, including DepthSplat [79], our reimplementation of GS-LRM [92], and our method pretrained solely on the first-stage static reconstruction.
+As shown in Table 2, although MoVieS is primarily designed for dynamic scenes, it maintains competitive performance on static scenes.
+Notably, when processing static inputs, our predicted motion naturally converges to zero (less than 1e-3), demonstrating the ability of MoVieS to implicitly differentiate between static and dynamic regions.
+4.2.2 Dynamic Scene View Synthesis
+We compare against three state-of-the-art open-sourced methods [72, 62, 26] for dynamic 3D Gaussian reconstruction, and also report static feed-forward baselines for reference.
+For DyCheck, we sample 13 frames of size $518\times 518$ from a random clip of 65 frames from iPhone-record videos, and include the other two cameras for evaluation.
+For NVIDIA, 12 input views at $379\times 672$ resolution are sampled in a round-robin style [13, 38, 33], where the i-th frame comes from the i-th camera at timestep i.
+As shown in Table 2, MoVieS achieves competitive or superior performance compared to baselines, while requiring only 0.93s per scene, which is orders of magnitude faster than prior approaches that rely on heavy pretrained models and complex multi-stage pipelines.
+Qualitative visualizations in Figure 3 further highlight the strength of our approach.
+While MoSca [26] demonstrates impressive performance, it struggles with sparse inputs and often overfits to seen poses, producing spiky and over-saturated artifacts under novel views and timesteps.
+In contrast, MoVieS leverages large-scale learned priors to generalize more effectively, yielding smoother and more realistic results.
+To ensure fair comparison and better reflect real scenarios, no video masks for dynamic objects are used in our experiments.
+It poses a major challenge for optimization-based methods like Shape-of-Motion [72] and Splatter-a-Video [62], which rely heavily on explicit motion segmentation.
+The challenge is particularly evident on the NVIDIA dataset, where significant camera shake hinders disentangling scene dynamics, leading to notably degraded performance and even worse than static baselines.
+In contrast, MoVieS exhibits strong robustness by directly learning to model motion.
+To evaluate the robustness of our method, visualization results on the in-the-wild DAVIS dataset [49] are provided in the supplementary material, where camera poses are estimated by MegaSaM [31].
+4.3 3D Point Tracking
+Trained on large-scale point tracking datasets [96, 21, 20], the proposed method can also densely track any 3D point corresponding to a pixel across video frames.
+We compare MoVieS against three strong baselines, including two state-of-the-art 2D point tracking methods, BootsTAP [10] and CoTracker3 [22], as well as a native 3D point tracking approach, SpatialTracker [77].
+For 2D trackers, a recent video depth estimation model [6] and ground-truth camera intrinsics are used to unproject tracked points into 3D space.
+Considering scale differences across methods, we normalize all predicted 3D points by their median norm before evaluation.
+Quantitative results are reported in Table 3.
+While 3D-based SpatialTracker generally outperforms 2D-based approaches, all of them rely heavily on pretrained monocular depth estimators for geometry reasoning, introducing significant noise and inconsistency in the 3D space.
+In contrast, MoVieS directly estimates 3D point positions in a shared world coordinate, enabling more accurate and robust 3D tracking, and achieves consistently superior or competitive performance across all datasets.
+Visualization results are provided in the supplementary material.
+4.4 Ablation and Analysis
 Camera Conditioning
-We investigate different camera conditioning strategies in the static pertaining stage.
-Quantitative comparisons are provided in Table 5.
-Camera tokens are injected throughout the feature backbone, enabling effective camera-aware modeling.
+We investigate different camera conditioning strategies in the static pretraining stage.
+As shown in Table 5, camera tokens are injected throughout the feature backbone, enabling effective camera-aware modeling.
 In contrast, Plücker embeddings provide limited conditioning on their own and are merely comparable to having no camera information at all.
@@
 Motion Supervision
-To learn object 3D movements in dynamic scenes, we provide two kinds of motion supervision: (1) point-wise L1 loss and (2) distribution loss, as described in Equation 4.
+To learn 3D movements in dynamic scenes, we provide two kinds of motion supervision: (1) point-wise L1 loss and (2) distribution loss, as described in Equation 5.
 Their efficiency is evaluated via the 3D point tracking task in Table 5.
 Without any motion supervision, i.e., learning solely from novel view synthesis, training exhibits severe loss oscillations and frequent None gradients.
-The distribution loss captures only relative motion between pixels, while point-wise L1 loss produces more reasonable motion maps.
+The distribution loss captures only relative motion between pixels, while the point-wise L1 loss produces more reasonable motion maps.
 Combining both leads to sharper boundaries.
@@
 Moreover, the depth head must now model both geometry and dynamics, increasing its burden and negatively affecting NVS.
-These results highlight the mutual reinforcement between NVS and motion estimation in MoVieS, where joint training leads to better performance on both.
+These results highlight the mutual reinforcement between NVS and motion estimation, where joint training leads to better performance on both.
+VGGT Initialization
+Although MoVieS is initialized by a pretrained model, i.e., VGGT [68], it is not an essential component of the proposed method.
+We verified it in our preliminary experiments that using a feature backbone trained entirely from scratch still leads to comparable performance with about 3 times slower convergence.
+The pretrained VGGT backbone simply serves to accelerate training, but does not fundamentally change the outcome.
+Specifically, we conduct quantitative experiments on DyCheck at $224\times 224$ resolution.
+MoVieS initialized from VGGT reaches 17.97 dB mPSNR within 2 days, whereas the counterpart trained from scratch achieves 16.64 dB after 6 days and has not yet fully converged, suggesting that continuous training could further improve its performance.
 4.5 Zero-shot Applications
@@
 Visualization results in Figure 5(a) present sharp edges and accurate motion directions.
-More results are provided in Figure 8.
+Different colors in the flow maps indicate different movement directions that are marked by block arrows.
+More visualization results are provided in the supplementary material.
 Moving Object Segmentation
-By thresholding the norm of per-pixel motion vectors, the estimated motion maps can be used to segment moving objects (Figure 5(b)), which is an essential task in computer vision and robotics (xie2024moving, ; huang2025segment, ).
-Remarkably, this is achieved without any explicit mask supervision, demonstrating the strong potential of our method.
-More results are provided in Figure 9.
-Figure 5: Zero-shot Applications.
-The predicted pixel-aligned motion maps from our model can be directly applied to downstream tasks, such as (a) scene flow estimation and (b) moving object segmentation, in a zero-shot manner, without any task-specific fine-tuning or supervision.
+By thresholding the norm of per-pixel motion vectors, the estimated motion maps can also be used to segment moving objects (Figure 5(b)), which is an essential task in computer vision and robotics [78, 18].
+Remarkably, this is achieved without any explicit supervision during training, demonstrating the strong potential of our method.
+More visualization results are provided in the supplementary material.
 5 Conclusion
-In this work, we introduce MoVieS, a feed-forward model for dynamic novel view synthesis from monocular videos.
-It’s trained on large-scale datasets from diverse sources and jointly models scene appearance, geometry and motion in a unified and efficient network.
+We introduce MoVieS, a feed-forward model for dynamic novel view synthesis from monocular videos.
+It’s trained on large-scale, diverse datasets and jointly models scene appearance, geometry, and motion in a unified and efficient framework.
 Dynamic splatter pixels are proposed to represent dynamic scenes, enabling accurate and temporally coherent 4D reconstruction.
-Beyond novel view synthesis, MoVieS supports a broad range of applications including depth estimation, 3D point tracking, scene flow estimation and moving object segmentation, demonstrating its versatility for dynamic scene perception.
-We hope this work could serve as a step toward generalizable dynamic scene understanding, and support applications that require spatial and motion intelligence.
-More discussions on limitations, future work and broader impact are included in Appendix B.
+MoVieS also supports various applications, such as depth estimation, 3D point tracking, scene flow estimation, and moving object segmentation, showcasing its versatility for dynamic scene perception.
+We hope this work could serve as a step toward generalizable dynamic scene understanding, and broader applications requiring spatial intelligence.
+Limitations and future work are discussed in the supplementary material.
+Supplementary Material
 Appendix A Implementation Details
 A.1 Dataset Details
-As summarized in Table 1, the majority of training datasets used in our experiments are synthetic, providing rich annotations such as pixel-aligned depth maps and accurate camera intrinsics and extrinsics.
+Table S.1: Training Datasets. Eight datasets from diverse sources are utilized to train MoVieS at scale. “#Repeat” denotes dataset duplication count during integration to balance their contributions.
+Dataset Dynamic? Depth? Tracking? Real? #Scenes #Frames #Repeat
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Detailed information about the training datasets used in this work is provided in Table S.1.
+The majority of training datasets used in our experiments are synthetic, providing rich annotations such as pixel-aligned depth maps and accurate camera intrinsics and extrinsics.
 Ground-truth 3D point trajectories are also available in PointOdyssey [96] and DynamicReplica [21].
-Stereo4D [35], on the other hand, is a large-scale dataset constructed from YouTube stereo videos and annotated using pretrained foundation models [62, 98].
+On the other hand, Stereo4D [20] is a large-scale dataset constructed from YouTube stereo videos and annotated using pretrained foundation models [64, 10].
 Despite being partially noisy, its diversity and scale offer strong generalization benefits.
-The utilized backbone VGGT [10] requires all 3D scenes to be normalized to a unit scale, which in turn necessitates depth maps and camera extrinsics to be defined in a consistent metric space.
-This requirement is satisfied by the synthetic datasets and Stereo4D, whose camera poses and depths are metric-aligned by construction.
-In contrast, RealEstate10K [28] only provides relative camera parameters estimated via COLMAP [6], resulting in an unknown global scale.
-To address this, we re-estimate both depth maps and camera extrinsics using recent foundation models, Video Depth Anything [5] and MegaSaM [101], to recover aligned geometry across frames.
+The pretrained backbone, VGGT [68], requires all 3D scenes to be normalized to a unit scale on average, which in turn necessitates depth maps and camera extrinsics to be defined in a consistent metric space.
+It is satisfied by all synthetic datasets and Stereo4D, whose camera poses and depths are metric-aligned by construction.
+However, RealEstate10K [97] only provides relative camera parameters estimated via COLMAP [56], resulting in an unknown global scale.
+To address this issue, we re-estimate both depth maps and camera extrinsics using recent foundation models, including Video Depth Anything [6], Depth Pro [2] and MegaSaM [31], to recover aligned geometry across frames.
 A.2 Curriculum Training
 We observed that training MoVieS is particularly unstable: the training loss often fluctuates abruptly, and gradients are prone to becoming None.
-It may arise from sparse annotations and the heterogeneous nature of the training data, which mixes datasets from various sources with differing domains (e.g., indoor vs.
+It may arise from sparse annotations and the heterogeneous nature of training datasets, which mix datasets from various sources with differing domains (e.g., indoor vs.
 outdoor, real vs.
-synthetic), camera fields of view, recording frame rates, etc.
+synthetic), camera FoV, recording frame rates, etc.
 Thanks to the versatile design of MoVieS, we employ a curriculum strategy that gradually increases the training complexity.
@@
 Since modeling dynamic scenes requires reconstructing a set of 3DGS for each query time, which results in high GPU memory usage, we start by training on 5 input views for dynamic scenes and then expand to 13 views for fine-tuning.
-Finally, the training resolution is increased to $518\times 518$.
+Finally, the training resolution is increased to $518$.
+Similar to VGGT [68], in the last training stage, we randomly sample the frame number from $2\sim 13$ and the aspect ratio from $0.5\sim 2$, with the largest side fixed to $518$.
 A.3 Training Details
-Image encoder, feature backbone and depth head of MoVieS are initialized from a geometrically pretrained transformer, VGGT [10].
-Motion head is initialized from its point head.
+Image encoder, feature backbone, and depth head of MoVieS are initialized from VGGT [68].
+Motion head is initialized from its pointmap head.
 Remaining components, such as the splatter head and camera/time embeddings, are trained from scratch.
@@
 MoVieS is then finetuned on $518\times 518$ videos with 13 frames using a learning rate of 1e-5, which takes around 1 day.
-To improve memory and computation efficiency, several techniques such as gsplat [93] rendering backend, DeepSpeed [94], gradient checkpointing [95], gradient accumulation, and bf16 mixed precision are adopted in this work.
-Appendix B Discussion
-B.1 Limitations and Future Work
+To improve memory and computation efficiency, several techniques such as gsplat [89] rendering backend, DeepSpeed [55], gradient checkpointing [7], gradient accumulation, and bf16 mixed precision are also adopted.
+For the multi-task training objective, we did not manually tune the relative weights between different loss terms.
+Instead, the weights were set to bring the numerical values of the losses into roughly similar ranges: $\lambda_{\text{d}}=\lambda_{\text{r}}=\lambda_{\text{pt}}=1$ and $\lambda_{\text{m}}=\lambda_{\text{dist}}=10$.
+Appendix B Limitations and Future Work
 Although MoVieS achieves competitive performance with inference speeds that are orders of magnitude faster than optimization-based methods, there remains a noticeable gap in reconstruction quality.
@@
 Seamlessly integrating camera pose estimation within the MoVieS pipeline could simplify the overall workflow, reduce error accumulation, and enhance adaptability to diverse scenarios.
-Moreover, scaling MoVieS to handle large-scale scenes and achieve high-resolution rendering remains a challenge.
+Moreover, scaling MoVieS to handle long videos and achieve high-resolution rendering remains a challenge.
 The computational cost and memory demand grow significantly with scene complexity and resolution, which constrains practical deployment.
@@
 Addressing these challenges will not only bridge the performance gap but also unlock the full potential of fast and high-quality dynamic scene reconstruction.
-B.2 Broader Impact
+Appendix C License Information
+We employ several open-source implementations in our experimental comparisons, including: (1) DepthSplat [79] [^1] https://github.com/cvg/depthsplat/tree/main (MIT License), (2) Splatter-a-Video [62] [^2] https://github.com/SunYangtian/Splatter_A_Video (Apache License), (3) Shape-of-Motion [72] [^3] https://github.com/vye16/shape-of-motion/ (MIT License), (4) MoSca [26] [^4] https://github.com/JiahuiLei/MoSca (MIT License), (5) BootsTAPIR [10] [^5] https://github.com/google-deepmind/tapnet (Apache License), (6) CoTracker3 [22] [^6] https://github.com/facebookresearch/co-tracker (Creative Commons Attribution-NonCommercial 4.0 International Public License), and (7) SpatialTracker [30] [^7] https://github.com/henry123-boy/SpaTracker (Attribution-NonCommercial 4.0 International).
+Datasets from diverse sources are utilized to train MoVieS, including: (1) RealEstate10K [97] [^8] https://google.github.io/realestate10k/ (Creative Commons Attribution 4.0 International License), (2) TartanAir [75] [^9] https://theairlab.org/tartanair-dataset/ (Creative Commons Attribution 4.0 International License), (3) MatrixCity [29] [^10] https://city-super.github.io/matrixcity/ (Apache License), (4) PointOdyssey [96] [^11] https://pointodyssey.com/ (MIT License), (5) DynamicReplica [21] [^12] https://github.com/facebookresearch/dynamic_stereo (Attribution-NonCommercial 4.0 International License), (6) Spring [44] [^13] https://spring-benchmark.org/ (CC BY 4.0), (7) VKITTI2 [3] [^14] https://europe.naverlabs.com/research/computer-vision/proxy-virtual-worlds-vkitti-2/ (Creative Commons Attribution-NonCommercial-ShareAlike 3.0), and (8) Stereo4D [20] [^15] https://stereo4d.github.io/ (CC0 1.0 Universal).
+Appendix D Broader Impact
 MoVieS provides substantial advantages for fields including robotics simulation, AR/VR, autonomous driving, and digital twins by enabling fast and generalizable dynamic scene understanding.
-Its ability to reconstruct dynamic environments efficiently can accelerate innovation and improve system performance in these areas.
+Its ability to efficiently reconstruct dynamic environments can accelerate innovation and improve system performance in these areas.
 However, such powerful technology also raises risks related to the unauthorized generation of content and potential privacy violations.
-To mitigate these risks, it is essential to establish clear ethical guidelines and implement appropriate regulatory measures to ensure the responsible and safe use of this technology.
-Appendix C More Visualization Results
-We provide more visualization results of MoVieS in Figure 6, 7, 8 and 9.
-Appendix D License Information
-We employ several open-source implementations in our experimental comparisons, including: (1) DepthSplat [24] [^1] https://github.com/cvg/depthsplat/tree/main (MIT License), (2) Splatter-a-Video [81] [^2] https://github.com/SunYangtian/Splatter_A_Video (Apache License), (3) Shape-of-Motion [80] [^3] https://github.com/vye16/shape-of-motion/ (MIT License), (4) MoSca [83] [^4] https://github.com/JiahuiLei/MoSca (MIT License), (5) BootsTAPIR [98] [^5] https://github.com/google-deepmind/tapnet (Apache License), (6) CoTracker3 [19] [^6] https://github.com/facebookresearch/co-tracker (Creative Commons Attribution-NonCommercial 4.0 International Public License), and (7) SpatialTracker [76] [^7] https://github.com/henry123-boy/SpaTracker (Attribution-NonCommercial 4.0 International).
-Several public datasets from diverse sources are utilized to train our models, including: (1) RealEstate10K [28] [^8] https://google.github.io/realestate10k/ (Creative Commons Attribution 4.0 International License), (2) TartanAir [29] [^9] https://theairlab.org/tartanair-dataset/ (Creative Commons Attribution 4.0 International License), (3) MatrixCity [30] [^10] https://city-super.github.io/matrixcity/ (Apache License), (4) PointOdyssey [33] [^11] https://pointodyssey.com/ (MIT License), (5) DynamicReplica [34] [^12] https://github.com/facebookresearch/dynamic_stereo (Attribution-NonCommercial 4.0 International License), (6) Spring [31] [^13] https://spring-benchmark.org/ (CC BY 4.0), (7) VKITTI2 [32] [^14] https://europe.naverlabs.com/research/computer-vision/proxy-virtual-worlds-vkitti-2/ (Creative Commons Attribution-NonCommercial-ShareAlike 3.0), and (8) Stereo4D [35] [^15] https://stereo4d.github.io/ (CC0 1.0 Universal).
-Figure 6: Qualitative results for reconstruction. “Motion” means the 3D points’ movements of the input frame with respect to the first one.
-Figure 7: Qualitative results for 3D point tracking. “Motion” means the 3D points’ movements of the input frame with respect to the first one.
-Figure 8: More results for scene flow estimation. “Flow” means the optical flow of the input pixels with respect to the first frame. Arrows on pictures roughly mark directions.
-Figure 9: More results for moving object segmentation. Masks for moving parts, which are filtered from motion maps, are highlighted across frames in light red.
+To mitigate these risks, it is essential to establish clear ethical guidelines and implement appropriate regulatory measures to ensure responsible and safe usage.
+Appendix E More Visualization Results
+We provide more visualization results on novel view synthesis, 3D point tracking, scene flow estimation, and dynamic object segmentation in Figure S.1, S.2, S.3 and S.4 respectively.
+Qualitative visualizations of novel view synthesis and 3D point tracking are conducted on the DAVIS dataset [49] and TAPVid-3D [25] respectively, which are not included in the training datasets.
+Other visualization results are conducted on the test set of the curated datasets to evaluate the generalizability of MoVieS.
+XYZ values in the 3D space of motion maps are normalized to $[0,1]$ and treated as RGB channels for visualization.
+Figure S.1: Qualitative results of novel view synthesis on the DAVIS dataset [49].
+Camera poses are estimated by MegaSaM [31].
+“Motion” denotes the 3D pixel displacements between the “Target Time” frame with respect to the “Target Viewpoint” frame.
+The “Novel View Synthesis” results show the reconstructed scene at the target time from the target viewpoint.
+Figure S.2: Qualitative results for 3D point tracking.
+“Motion” means the 3D points’ movements of the input frame with respect to the first one.
+Points in the 3D space are projected to the image space for 2D point tracking visualization.
+Figure S.3: Qualitative results for scene flow estimation.
+“Flow” means the optical flow of the input pixels with respect to the first frame, and is obtained by projecting the estimated motion maps in the 3D space to the camera space.
+Each color of optical flow means a specific 2D direction and arrows on pictures roughly mark the directions.
+Figure S.4: Qualitative results for moving object segmentation.
+Masks for moving parts, which are filtered from the values of estimated motion maps, are highlighted across frames in light red.
+Regions with high motion values are regarded as moving parts.